### v1

# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

Abstract
Large Language Models (LLMs) are increasingly being applied to specialized, high-stakes domains like engineering, which demands rigorous evaluation of their complex reasoning capabilities. While current benchmarks assess language understanding, factual recall, mathematics or code generation, none capture the integrative reasoning central to engineering where scientific principles, quantitative modeling and practical constraints must converge. To address this gap, we introduce EngChain, a benchmark for verifiable multi-step engineering problem-solving. EngChain contains 90 problems spanning three engineering branches, organized into 9 domains and 20 distinct areas. The problems are generated from symbolic templates with a high degree of randomization to ensure diversity and eliminate the risk of contamination. With this benchmark, we move beyond final answer accuracy with a two-stage evaluation: we first quantitatively verify the numerical and semantic validity of each reasoning step and then introduce LLM-as-a-Judge, an automated system to qualitatively categorize the identified reasoning errors.
EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering
Ayesha Gull1 Muhammad Usman Safder1 Rania Elbadry2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI ProjectCodeApp
**footnotetext: Equal contribution.
1 Introduction
As large language models (LLMs) transition from general-purpose language tasks to specialized, high-stakes domains such as scientific discovery, industrial design, and engineering analysis, the need for rigorous and meaningful evaluation has become significant (Zhao et al., 2023; Xie et al., 2023a). The progress of the field is gated by the quality of its benchmarks: without tests that capture practical, physically grounded reasoning, it is difficult to separate genuine capability from sophisticated mimicry.
To meet this challenge, we introduce EngChain, a new benchmark for verifiable, multi-step engineering problem-solving. Our benchmark is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique, verifiable problems. Figure 1 provides a concrete example from our benchmark, illustrating how a single template generates a complete problem instance along with its full chain-of-thought solution.
Figure 1: An example from EngChain (CSTR Volume Calculation). A symbolic template (1) generates a unique problem instance (2) and its verifiable, step-by-step chain-of-thought solution (3).
This design directly targets two key problems in current LLM evaluation. The first is “benchmark saturation” (Ott et al., 2022), where models rapidly achieve superhuman performance on static benchmarks like GLUE (Wang et al., 2019) and even their more complex successors such as BBH (Kazemi et al., 2025). The second is a structural limitation: most benchmarks evaluate critical skills in disciplinary silos. General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., MATH (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively.
This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012). This holistic, procedural reasoning is not captured by existing benchmarks (e.g., Felten et al., 2025; Li et al., 2025b), which predominantly focus on final-answer accuracy. In engineering, where a flawed process can lead to catastrophic failure, this evaluation gap is critical. To address this gap, EngChain uses symbolic templates to produce integrative reasoning problems and a two-stage evaluation via LLM-as-a-Judge that verifies the full reasoning chain rather than just the final answer.
Verifiable multi-step reasoning is fundamental to building trustworthy AI systems across all domains. Concurrent with our work, FinChain (Xie et al., 2025) introduced a symbolic benchmark for verifiable chain-of-thought financial reasoning. However, the two benchmarks address fundamentally different reasoning types. FinChain evaluates reasoning based on economic principles, whereas EngChain evaluates reasoning grounded in the physical laws and practical constraints of engineering. Moreover, FinChain uses a global alignment metric to score the entire reasoning path, while EngChain employs a two-stage evaluation that provides both a quantitative Step F1-Score for procedural correctness and a qualitative, diagnostic error analysis using our LLM-as-a-Judge system.
In summary, our contributions are threefold:
  1. 1.
    We introduce EngChain, a new symbolic benchmark for verifiable, multi-step engineering problem-solving with 90 problem templates across nine domains.
  2. 2.
    We propose a rigorous two-stage evaluation framework, including our qualitative LLM-as-a-Judge system, that moves beyond final answer accuracy to diagnose reasoning failures.
  3. 3.
    We present a comprehensive analysis of 11 frontier LLMs, revealing a critical “right for the wrong reasons" gap and identifying that the majority of reasoning failures are due to Conceptual Errors, not calculation slips.
2 Related Work
2.1 Benchmarks in Mathematics and Coding
A large body of work on multi-step reasoning benchmarks has focused on foundational, yet isolated, cognitive skills, particularly in mathematics and coding. In mathematical reasoning, benchmarks ranging from grade-school problems (GSM8K)  (Cobbe et al., 2021) to high-school and college-level competition problems (MATH)  (Hendrycks et al., 2021b) establish a clear standard for evaluating abstract logical deduction. This is further refined by HARDMATH  (Fan et al., 2024) for applied mathematics and GSM-Symbolic  (Mirzadeh et al., 2024) for programming and symbolic tasks. Similarly, benchmarks for coding and software engineering evaluate models across different levels of complexity. HumanEval  (Chen et al., 2021) focuses on translating natural language instructions into small, correct code snippets. MBPP  (Austin et al., 2021) and its derivative MBUPP extend this by testing models on a broader set of programming problems with varying difficulty. SWE-Bench  (Jimenez et al., 2024) addresses complex, real-world codebases, measuring a model’s ability to solve practical software engineering challenges. While these benchmarks are essential for capturing foundational skills, their focus on isolated abilities limits their usefulness for engineering. A model can excel at solving abstract equations or writing pure algorithms yet still be unable to apply those skills to a problem constrained by physical laws and material properties.
2.2 Benchmarks in the Physical Sciences
Recognizing this limitation, later benchmarks focus on applied reasoning within the physical sciences. UGPhysics  (Xu et al., 2025) provides broad coverage and tests how models apply mathematical reasoning to physical laws. Building on this, PhysReason  (Zhang et al., 2025a) emphasizes process-oriented problem solving, guiding models through step-by-step solutions. ABench-Physics  (Zhang et al., 2025b) introduces highly challenging tasks to evaluate advanced capabilities, while PHYBench  (Qiu et al., 2025) targets symbolic computation. NEWTON  (Wang et al., 2023) assesses qualitative understanding, complementing the more quantitative approaches. Other benchmarks, such as LLM-SRBench  (Shojaee et al., 2025), focus specifically on quantitative scientific modeling. Together, these benchmarks show how models apply mathematical reasoning to well-defined physical laws, marking an important advance towards assessing real-world problem-solving. However, they remain largely confined to physics and have not been widely adopted for evaluating models in real-world engineering practices.
2.3 Benchmarks in Engineering
Current approaches to evaluating engineering reasoning are fractured, split between benchmarks that are overly broad and those that are deep but methodologically limited. On one hand, generalist benchmarks like MMLU  (Hendrycks et al., 2021a) and BIG-Bench  (Luo et al., 2024) include engineering topics, but their reliance on multiple-choice, knowledge-based questions primarily tests factual recall rather than the deep, procedural reasoning required for practical engineering problem-solving. On the other hand, more specialized engineering benchmarks, while deeper in scope, either remain confined to a single scientific discipline or lack a scalable, deterministic method for verifying the reasoning process itself. For example, EngiBench  (Felten et al., 2025) provides broad coverage but uses rubric-based scoring for its more complex problems, whereas FEA-Bench  (Li et al., 2025b) evaluates a model’s ability to use professional simulation software, yet this measures tool proficiency rather than fundamental reasoning.
The preceding review makes it clear that the field lacks a benchmark that captures the integrative, cross-disciplinary, and verifiable reasoning central to engineering. To address this critical gap, we introduce EngChain, a benchmark for verifiable, multi-step engineering reasoning. It uses a two-stage evaluation that combines quantitative verification of each intermediate step with a qualitative error diagnosis, covering 90 distinct problem types across nine domains and three engineering branches.
3 EngChain Taxonomy
EngChain covers a wide range of core engineering problems, focusing on domains that are central to real-world practice, which we show in Figure 2.
In Chemical Engineering, it targets the three pillars of chemical process design and analysis. Kinetics problems test the modeling of reaction rates and mechanisms, requiring reasoning about how systems evolve over time. Thermodynamics problems center on energy balances and phase behavior, capturing the constraints that govern feasibility and efficiency. Transport phenomena problems address momentum, heat, and mass transfer, linking microscopic principles to large-scale process behavior. In Electrical Engineering, it covers the full spectrum from fundamental physics to applied systems. At the foundation, it includes electromagnetics, which governs the behavior of fields and waves. Building on this, it addresses signals and systems, focusing on representation, transformation, and filtering of information. Finally, it extends to communications, where these principles are applied to the reliable transmission and processing of signals in real-world contexts. In Mechanical Engineering, it spans three core domains. Fluid mechanics problems address laminar and turbulent flow, testing how models capture momentum and energy transfer in fluids. Mechanics of materials emphasizes stress–strain relationships, deformation, and failure analysis, reflecting how structures respond to applied loads. Vibrations problems focus on natural frequencies and damping, capturing the dynamic stability of mechanical systems. Together, these areas test how forces, motion, and stability interact in physical systems.
Figure 2: The taxonomy of EngChain.
By integrating these principle subfields across the main branches of engineering, EngChain provides a comprehensive foundation of the analytical and procedural reasoning skills essential for engineering problem-solving.
4 EngChain Design and Methodology
To ensure the rigor, scalability, and long-term viability of the benchmark, we adopted a programmatic, template-based methodology. Unlike static datasets, which are finite and susceptible to being inadvertently included in the training corpora of future models, a programmatic approach offers several key advantages. It provides virtually limitless scalability and diversity, as a single template can generate thousands of unique problem instances. This design also ensures strong resistance to training data contamination, as it tests for generalizable reasoning on novel problems rather than the memorization of previously seen solutions.
4.1 Content Selection Framework
The selection of content for EngChain followed a principled, multi-stage methodology designed to ensure the benchmark is both representative of a formal engineering curriculum and grounded in foundational, real-world principles. The process involved a hierarchical selection of domains, areas, and finally, the specific problems to be templatized.
4.1.1 Domain and Area Selection
The nine core domains that form the backbone of EngChain were identified using a two-pronged approach. First, a thorough literature review was conducted, surveying established engineering taxonomies and curricular standards set by leading professional and accreditation bodies. This involved analyzing the program criteria from organizations like the Accreditation Board for Engineering and Technology (ABET, 2024), which mandates core competencies in areas like thermodynamics and circuit analysis for accredited degrees. We also reviewed the educational frameworks of major professional societies, such as the American Society of Mechanical Engineers (ASME, 2025), the Institute of Electrical and Electronics Engineers (IEEE, 2025), and the American Institute of Chemical Engineers (AIChE, 2025), to ensure our selected domains represent a consensus on the foundational knowledge required for professional practice.
Second, these findings were cross-referenced by querying an LLM (gemini-2.0-flash). The model was prompted to identify and rank the most fundamental, cornerstone domains within each of the three primary engineering branches. The full prompt used for this domain-level validation is provided in Appendix A. The final set of nine domains represents a strong consensus between traditional academic structures and the knowledge synthesized by modern AI. A similar two-pronged methodology was then used to identify the most critical areas within each domain: first, by extracting the core pedagogical units from the authoritative textbooks, and second, by cross-referencing this list using a LLM (gemini-2.0-flash). The full prompt used for this area-level validation is provided in Appendix B.
4.1.2 Template Selection
Once the domains and areas were defined, we developed the 90 problem templates. The core engineering principles, problem typologies, and typical constraints for these templates were sourced from the authoritative and widely-adopted textbooks in each field. This textbook-driven approach ensures that every problem in EngChain is pedagogically sound and central to engineering practice. A complete list of these foundational texts is provided in Appendix C.
Crucially, the number of templates developed for each area was not uniform. Instead, it was guided by a Pedagogical Significance Score assigned to each area. This score was programmatically determined by querying an expert-persona LLM (gemini-2.0-flash) to rate each area’s centrality to the curriculum on a 1-5 scale, as detailed in Appendix D. Foundational concepts with higher scores received more extensive coverage. For example, within the Reaction Kinetics domain, the area of Mole Balances (Significance Score: 5) was covered by five templates, while the more specialized application of Levenspiel Plot Interpretation (Significance Score: 1) was covered by a single template. Therefore, the varying number of templates per area is a direct and intentional outcome of our methodology. This ensures that EngChain’s structure reflects the hierarchical and foundational nature of engineering knowledge as it is taught and practiced.
4.2 Template Generation Pipeline
The template generation process follows a systematic pipeline as shown in Figure 3. First, we generate domain-aware parameters are generated. Next, the core engineering computations are performed. These results are then thoroughly validated to ensure they are physically plausible. Finally, the validated content is assembled into a natural language question and its corresponding step-by-step solution.
Figure 3: Our template generation pipeline.
We format the output strings to be both human-readable, like a textbook example, and machine-parsable for automated evaluation. After template functions are defined, we generate the problem instances by repeatedly calling each template, with every run using a unique random seed for reproducibility.
4.3 Template Structure and Outputs
The core of our benchmark is the template function, a standardized Python function that serves as the blueprint for a specific type of engineering problem. To ensure consistency, every template sticks to a simple, standardized signature i.e., it takes no arguments and returns a (question, solution) tuple. The question is a formatted string containing the problem statement to be presented to the LLM, whereas the solution is a formatted string containing the complete reasoning trace rather than only a numerical answer. A generic multi-disciplinary template structure is provided in Figure 4.
Figure 4: Generic multi-disciplinary template structure.
Each solution is structured like a textbook example such that the principles are stated, values are substituted, and intermediate steps are shown with consistent units.
4.4 Domain-Aware Parameterization and Sanity Checks
EngChain uses domain-aware parameterization, ensuring that problems are not just mathematically solvable but are also grounded in physical and engineering realism. This is done by constraining all generated values by the principles of the domain being tested.
For instance, in Chemical Engineering, problems use real reactants like Benzene, pre-validated stoichiometric equations, and temperature dependent heat capacity parameters rather than random constant values. In Electrical Engineering, parameters are grounded in fundamental constants (e.g., $c_{0}$, $\epsilon_{0}$) and real-world materials (e.g., Polyethylene), with signal properties procedurally generated to explicitly test concepts like aliasing. In Mechanical Engineering, problems co-select consistent properties for real materials (e.g., 6061-T6 Aluminum) and are context-aware, such as using Mercury for manometer problems. A comprehensive list of these parameterization examples is provided in Appendix E, and the authoritative data sources used are listed in Appendix F.
Furthermore, across all templates, these already constrained parameters are subjected to multi-layer sanity checks. For example, a chemical engineering template verifies that a final concentration is less than the initial one. A mechanical engineering problem ensures a calculated stress does not exceed the material’s yield strength. Similarly, an electrical engineering template confirms a wave’s propagation speed does not exceed the speed of light. These checks validate that the final combination of values results in a solvable and physically meaningful scenario.
4.5 AI-Assisted Quality Assurance
To ensure the correctness and scalability of the benchmark, we developed an AI-Assisted Quality Assurance workflow. While traditional unit tests are effective, they proved impractical for our generative framework due to the high number of templates and the complexity of verifying physically plausible outputs. Our quality assurance workflow automates this process by employing an LLM as a first-pass Peer Reviewer to validate new problem templates before their inclusion in the final benchmark.
The core of this workflow is a programmatic validation script that prepares a comprehensive validation package for the LLM. For each template, this package includes not only the template’s full Python source code but also three unique instances (question-solution pairs) generated by it. This context is crucial, as it allows the LLM to assess both the static logic of the code and the dynamic quality of its output. We then pass this package to the LLM via a carefully engineered prompt. As detailed in Appendix G, the prompt instructs the LLM to assume the persona of an expert engineering professor and to evaluate the template against a multi-axis rubric provided in Appendix H.
Upon receiving the results, our system programmatically parses and evaluates the scores. A template, $T$, is automatically approved if and only if the score assigned to it for every core category ($c_{i}$) in our rubric meets or exceeds a predefined threshold, $\theta$. We formalize this approval condition as:
$\forall c_{i}\in C,\text{Score}(T,c_{i})\geq\theta$
We established a high threshold of $\theta=4$ based on empirical testing against a control set of known-good and deliberately flawed templates (Examples are provided in Appendix L). Any template failing to meet this criterion is automatically flagged for human review. This threshold demonstrated an optimal balance between maximizing automation and reliably catching potential errors.
In practice, this automated review process proved highly effective. Out of the 90 total templates, only 5 (5.55%) were flagged by the LLM and required subsequent human review. [^1] This human review was performed by the authors, who possess relevant engineering domain expertise. While we did not calculate formal inter-rater reliability between the LLM and human reviewers for this quality assurance step, we manually inspected and corrected the flagged templates.
4.6 Symbolic Generation and Combinatorial Diversity
A core feature of EngChain is its programmatic generation, designed to create a problem space so vast that it effectively prevents evaluation through memorization. Rather than relying on a static set of problems, EngChain is built upon symbolic templates where parameters are systematically randomized within domain-aware, physical constraints.
This design provides strong resistance to training data contamination, as it makes it statistically improbable that an LLM has encountered any specific problem instance. Consequently, successful performance on EngChain is stronger evidence of a model’s ability to generalize its reasoning capabilities. The model is forced to solve each problem from first principles rather than relying on pattern matching or rote recall of seen solutions.
4.7 Difficulty Scaling
Problem difficulty in EngChain is not a subjective label but a calibrated measure determined by systematically varying three primary factors:
  • Conceptual Complexity refers to the number and integration of engineering principles required. An Easy problem typically tests a single, isolated concept like Ohm’s Law. An Intermediate problem requires the combination of two or three concepts, such as applying the principle of superposition to calculate a net electric field from multiple point charges. An Advanced problem demands the integration of multiple complex ideas, such as using numerical integration on Levenspiel plot data to determine a reactor volume.
  • Mathematical Sophistication describes the nature of the required calculations. Easy problems involve direct algebraic substitution. Intermediate problems may require basic calculus or vector addition. Advanced problems escalate to solving differential equations or applying numerical methods, such as the trapezoidal rule for integration.
  • Procedural Depth measures the length and interdependence of the reasoning chain. An Easy problem might be solvable in one or two steps. In contrast, an Advanced problem involves a long sequence of calculations where the output of each step is a critical input for the next, testing a model’s ability to maintain a coherent and accurate computational trace.
This multi-axis framework for difficulty allows EngChain to move beyond a simple easy/hard dichotomy and provides a more fine-grained analysis of an LLM’s reasoning capabilities.
5 Evaluation Framework
For a verifiable assessment of the entire reasoning process, we compute several key metrics by comparing a model’s generated solution against the ground-truth reasoning trace. The evaluation is multifaceted. It jointly evaluates final answer accuracy, the fluency and semantic quality of the generated solution, and the correctness of the multi-step reasoning chain.
5.1 Final Answer Accuracy
It is determined by comparing the model’s predicted numerical answer ($A_{pred}$) to the ground-truth answer ($A_{gt}$). A response is considered correct only if its relative error is less than a strict tolerance ($\epsilon_{final}$) of 1%, a standard appropriate for engineering precision. The calculation is formalized as follows:
$\text{Relative Error}=\left|\frac{A_{gt}-A_{pred}}{A_{gt}}\right|<\epsilon_{final}$
5.2 Textual Quality
We assess the textual quality of a generated solution, evaluating its fluency and semantic fidelity against the ground-truth text using a combination of semantic and lexical similarity scores using the most common reference based evaluation metrics (Xie et al., 2023b):
  • BERTScore (Zhang et al., 2020) measures deep semantic similarity by aligning contextual embeddings of tokens from the generated solution and the ground-truth solution. It computes precision, recall, and an F1-score to provide a robust measure of semantic equivalence, capturing nuances that lexical overlap metrics may miss.
  • ROUGE-2 (Lin, 2004) quantifies the lexical overlap by calculating the F1-score based on the number of matching adjacent word pairs (bigrams) between the generated and ground-truth texts.
Together, these scores provide a complementary assessment that reflects the model’s ability to produce solutions using appropriate and correctly structured domain-specific terminology.
5.3 Multi-Step Reasoning Verification
This is the core of our evaluation framework which assesses the entire logical chain of a model’s solution. This process moves beyond final answer accuracy by comparing each predicted reasoning step ($P_{j}$) against each ground-truth step ($GT_{i}$). A step is considered valid only if it satisfies two criteria: semantic similarity and numerical correctness.
To formalize this, we compute a combined similarity matrix, $S_{comb}$, by taking the element-wise product (Hadamard product $\circ$) of a semantic similarity matrix ($S_{sem}$) and a numerical correctness matrix ($S_{num}$).
$S_{comb}=S_{sem}\circ S_{num}$
  • Semantic Similarity ($S_{sem}$) is computed using a powerful Cross-Encoder model (cross-encoder/stsb-roberta-large). This model processes each ground-truth and predicted step as a single contextual pair, yielding a high-fidelity similarity score that captures nuanced semantic relationships more effectively than bi-encoder models like the initially used SentenceTransformer.
  • Numerical Correctness ($S_{num}$) is a binary matrix where an entry is 1 if the intermediate numerical answer in a predicted step matches the corresponding ground-truth step within a 2% relative error tolerance, and 0 otherwise.
From this combined matrix, we determine if a step is “matched” by checking if its maximum similarity score exceeds a predefined threshold, $\tau$. We then calculate recall and precision to evaluate the reasoning chain. Recall measures the fraction of ground-truth steps the model successfully reproduced:
$\text{Recall}=\frac{1}{|GT|}\sum_{i=1}^{|GT|}\mathbb{I}\left(\max_{j}S_{comb}[i,j]>\tau\right)$
Precision measures the fraction of the model’s generated steps that were valid and relevant:
$\text{Precision}=\frac{1}{|P|}\sum_{j=1}^{|P|}\mathbb{I}\left(\max_{i}S_{comb}[i,j]>\tau\right)$
Finally, our primary metric for reasoning quality, the Step F1-Score, is calculated as the harmonic mean of the step-level recall and precision defined above, providing a single, robust measure of procedural correctness.
$\text{Step F1-Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
5.4 Automated Error Analysis with LLM-as-a-Judge
While the Step F1-Score provides a sound quantitative measure of a model’s procedural correctness, it has a significant limitation: it can identify that a reasoning chain is flawed, but it cannot explain why. For instance, as will be discussed later, our preliminary results revealed models that achieved high final-answer accuracy despite very low reasoning scores, a classic case of being “right for the wrong reasons.” This made us realize that to diagnose these failures and determine their cause, a deeper, more qualitative analysis was required. Hence, to move beyond quantitative scores and introduce a layer of explainability, we propose a novel Automated Error Analysis system that employs a large language model as an expert evaluator, or an LLM-as-a-Judge.
This system is triggered for any ground-truth step ($GT_{i}$) that fails to find a satisfactory match among the predicted steps. A step is flagged as a “failure” if its maximum combined similarity score falls below an empirically determined threshold, $\tau=0.7$. This condition is formalized as:
$\max_{j}S_{comb}[i,j]<\tau$
Figure 5: Automated error analysis with the LLM-as-a-Judge workflow.
The overall workflow of this automated analysis process is illustrated in Figure 5. At its core, the process is rooted in carefully designed prompt engineering, where the judge model is given a system prompt defining its persona as an expert engineering professor and a set of critical rules for logical consistency. The input to the judge is a structured prompt containing the full problem context, the ground-truth step, and the model’s incorrect step. Providing this complete context is crucial for the judge to accurately classify the failure into one of the five distinct categories, which are defined in Appendix I. To ensure a reliably parsable output, the judge is instructed to return only a structured JSON object containing its classification and a detailed and scalable diagnosis of the model’s specific failure modes. The full system prompt used for the automated error analysis is provided in Appendix J.
6 Experiments
6.1 Evaluation Model Suite
To provide a comprehensive evaluation of engineering reasoning capabilities, we evaluate models from a diverse suite. The results presented in Section 7 focus on 11 distinct models from the first category. While a full evaluation of the other categories is planned for future work, we group the complete suite of models into the following four-way taxonomy:
  1. 1.
    Frontier proprietary models, which represent the current upper bound of performance. This includes models such as GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025).
  2. 2.
    Science/Engineering-enhanced models, such as Galactica (Taylor et al., 2022), Minerva (Lewkowycz et al., 2022), ChemLLM (Zhang et al., 2024), SciLit01 (Li et al., 2025a), and SciAgent (Gao et al., 2024), are systems pre-trained or fine-tuned on vast scientific and technical corpora to improve domain-specific knowledge and complex procedural reasoning.
  3. 3.
    Math-enhanced models, such as WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024), are systems specifically trained on mathematical and symbolic corpora to enhance their quantitative reasoning capabilities.
  4. 4.
    General-purpose open models, such as LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025), provide strong, domain-agnostic foundations that are often used to create specialized variants.
This four-way taxonomy provides a structure for comparing different model architectures and training paradigms, highlighting how model scale, domain specialization, and mathematical supervision influence symbolic engineering reasoning. The detailed configurations and model sources can be found in Appendix K.
6.2 Experimental Setup
We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model. We evaluate all models under a unified decoding configuration: a low temperature of 0.2 and a maximum token limit of 4,096. We use a low temperature (T=0.2) to ensure deterministic, high-confidence outputs and a high token limit (4,096) to allow ample space for detailed, multi-step solutions without premature cut-offs. We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
You are an expert engineer. Solve the following
problem by providing a detailed, structured
solution. Use the exact headings and
formatting provided below.
# Given
List all known variables and their
values with units.
# Find
State the variable(s) to be calculated.
# Formulae
Write down all necessary governing equations
before substituting any values.
# Solution
Provide a step-by-step calculation. Each step
must start on a new line and be formatted exactly
as ’**Step X:**’, where X is the step number.
Show the substitution of values into
the formulae clearly.
# Final Answer
State the final numerical result with its units
in the format: **Answer:** [value] [units]
To process the model outputs, we first segment the text into reasoning steps. We use a regular expression to identify line-start markers that match either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.). A subsequent regex cleans these prefixes. We then extract numerical values from each step and the final answer using a three-priority heuristic: first, matching the **Answer:** tag; second, falling back to the last number following an equals sign (e.g., V = 100.0); and third, taking the last standalone number. If no overall **Answer:** tag is found, the final numerical value from the last step is used for verification.
7 Results
7.1 Overall Model Performance
The primary finding from our evaluation is a stark and universal divergence between the models’ ability to produce a correct final answer and the verifiable correctness of their reasoning process. As shown in  Table 1, while the top models achieve a final answer accuracy of approximately 63.1% (gpt-4.1-mini and gpt-5-mini), their procedural reasoning capability is critically low. The Reasoning (F1 Score), our primary metric for procedural correctness, is alarmingly poor across the board, with the top-performing model, gemini-2.5-flash, only achieving 19.32%. This demonstrates that in the vast majority of cases, all models fail to follow a sound, verifiable problem-solving methodology, confirming the “right for the wrong reasons” phenomenon at scale.
This significant reasoning gap is further masked by high textual fluency. The BERTScore, for instance, remained consistently high for all models, averaging approximately 0.87. This indicates that models excel at generating semantically appropriate, engineering-domain language, even when the underlying procedural logic is flawed or nonsensical.
Model Final Answer Accuracy $\uparrow$ Reasoning (F1 Score) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
Frontier Proprietary Models
Claude 3.7 56.8149.55 16.9121.76 88.031.16 29.287.65 37.097.63
Claude 4 57.7849.41 14.1618.28 87.931.33 29.558.16 37.257.91
Claude 4.5 52.2249.97 14.1919.72 87.391.09 28.108.15 35.347.58
Deepseek v3.2 61.2648.73 13.5319.54 86.541.23 25.698.96 34.579.12
Deepseek R1 59.9349.02 19.0324.82 86.481.34 25.588.34 34.498.30
Gemini 2.5 Flash 59.4149.13 19.3223.56 87.551.63 28.829.17 36.409.36
Gemini 2.5 Pro 59.9349.02 17.9721.01 88.241.41 31.828.29 37.918.49
GPT 4.1 57.7849.41 12.1315.85 86.651.24 26.749.13 34.459.12
GPT 4.1 Mini 63.1148.27 14.8419.69 86.521.14 25.198.17 33.828.34
GPT 5 60.4448.92 15.4721.79 88.191.21 28.988.30 36.558.06
GPT 5 Mini 63.1148.27 14.3020.32 87.951.10 26.637.62 34.747.52
Table 1: Overall performance of frontier models on the EngChain benchmark (N=1350) under a Zero-Shot setting. Scores represent the mean, with standard deviation shown in superscript. The highest mean score in each column is in bold.
7.2 Performance by Engineering Branch
A fine-grained analysis of the results reveals a significant performance disparity across the three engineering branches, as illustrated in Figure 6. We observe that model reasoning capabilities are not uniform and are highly sensitive to the specific scientific principles and mathematical structures of a given engineering domain.
Chemical Engineering consistently proved to be the most challenging domain for all models, yielding the lowest scores in both final answer accuracy and procedural reasoning. The reasoning collapse in this domain is particularly severe, with several models falling into single-digit Reasoning (F1 Score)s, including Deepseek v3.2 (6.90%) and GPT 5 Mini (7.58%).
Conversely, Mechanical Engineering was clearly the most solvable domain, where models achieved their highest scores. This branch produced the peak Final Answer Accuracy of the entire benchmark (Gemini 2.5 Pro at 70.22%) and the highest Reasoning (F1 Score) (Deepseek R1 at 30.11%).
Electrical Engineering demonstrated intermediate performance, with scores that consistently fell between the other two branches. Its peak Reasoning (F1 Score) of 22.49% (Gemini 2.5 Flash) and peak Final Answer Accuracy of 67.11% (GPT 4.1 Mini) were modest, clearly lagging behind the top scores in Mechanical Engineering but remaining well above the challenging baseline set by Chemical Engineering.
Figure 6: Branch-level performance comparison of the top four models ranked by mean Reasoning (F1 Score) across the three engineering branches. Chemical Engineering is consistently the most challenging domain for procedural reasoning.
7.3 Performance by Engineering Domain
Further analysis reveals that model performance is highly inconsistent across the nine distinct engineering domains, as visualized in the radar plot in Figure 7. This “spiky” performance profile, where a model can excel in one domain while failing in another, suggests that current models possess patches of specialized knowledge rather than generalized, first-principles reasoning.
The data shows several domains are exceptionally challenging for all models. Domains such as Reaction Kinetics, Mechanics of Materials, Transport Phenomena, and Electromagnetics & Waves produced critically low reasoning scores. In these areas, it was common for top-tier models to score in the single digits, with Deepseek v3.2 scoring 6.90% in Reaction Kinetics and GPT 5 Mini scoring 7.58% in Mechanics of Materials.
Conversely, the domain of Vibrations & Acoustics emerged as a clear outlier, representing the easiest domain for nearly all models. The reasoning scores here were dramatically higher than in any other domain, with Deepseek R1 achieving 52.26% and Gemini 2.5 Flash achieving 40.70%. The domain of Signals & Systems was another area of relative strength, where most models achieved some of their highest reasoning scores.
This domain-level variance underscores the need for broad, multi-domain benchmarks. We hypothesize this performance gap is linked to the underlying mathematical complexity. The high-scoring domains in our results often feature problems solvable with direct algebraic substitution or linear, discrete procedures. The most challenging domains, in contrast, frequently require models to set up and solve differential equations or manage systems of complex, interdependent variables, which appear to be critical failure points for current procedural reasoning.
Figure 7: Domain-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot reveals a “spiky” performance profile, highlighting that model reasoning is highly specialized and not generalized across topics.
7.4 Performance by Problem Difficulty
As expected, model performance degrades as problem complexity increases, but this trend is most pronounced in Final Answer Accuracy. Most models maintain a respectable accuracy of 60-70% on Easy and Intermediate tasks. However, performance drops sharply on Advanced problems, with nearly all models falling into the 40-50% range. For instance, gemini-2.5-pro’s accuracy falls from 68.6% on Intermediate problems to just 45.2% on Advanced.
The Reasoning (F1 Score), by contrast, tells a more complex story, as illustrated in Figure 8. Instead of a uniform decline, the reasoning scores for many top models (such as claude-3-7 and gpt-5) are lowest on Intermediate tasks before recovering slightly on Advanced tasks. A few models (e.g., gemini-2.5-flash and deepseek-reasoner) even show a slight improvement in reasoning as difficulty increases. The most critical finding, however, is that reasoning scores remain universally low (sub-21%) across all three difficulty levels. This suggests the models’ procedural reasoning failures are fundamental and not simply a function of task complexity.
Figure 8: Difficulty-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot shows procedural reasoning performance is critically low across all levels of complexity, not just on advanced tasks.
7.5 Qualitative Error Analysis with LLM-as-Judge
While the quantitative Reasoning (F1 Score) identifies that a reasoning chain is flawed, it cannot explain why. To diagnose the root causes of these failures, we employed our LLM-as-Judge system (described in Section 5) to perform a qualitative error analysis on Gemini 2.5 Pro. We selected this model as our representative subject because it is a strong “all-rounder,” achieving both a top-tier reasoning score (17.97%) and the first rank on all language fluency metrics (BERTScore, ROUGE-2, and ROUGE-L). The analysis was done using GPT-4o (OpenAI, 2024) as the judge or expert evaluator model.
The judge was used to analyze all 4,118 reasoning steps that were flagged by our verifier as a “mismatch” from the ground-truth solution. The results, summarized in Table 2, are striking. The single most significant finding is that the vast majority of flagged steps were not errors at all. Our judge classified 73.94% (3,045 instances) as Alternative Correct or in other words, valid reasoning paths that simply differed from the ground-truth solution. This overwhelmingly indicates that models are frequently penalized by rigid, single-path evaluations for finding a valid, creative, or different solution, a phenomenon masked by low quantitative scores.
For the remaining 1,072 instances identified as genuine errors, Conceptual Error was the dominant failure mode, accounting for 59.1% of all true errors (634 instances). This demonstrates that models are far more likely to fail because they misapply a physical principle or choose the wrong formula, rather than making simple Calculation Errors (14.3%) or Input Errors (7.6%). This finding pinpoints the primary weakness of modern LLMs not in their arithmetic, but in their application of deep, domain-specific knowledge. A full breakdown of error distributions by branch, domain, and difficulty can be found in Appendix M.
Error Category Total Count Percentage (%)
Alternative Correct 3,045 73.94%
Conceptual Error 634 15.40%
Other 204 4.95%
Calculation Error 153 3.72%
Input Error 81 1.97%
Analysis Failed 1 0.02%
Total 4,118 100.0%
Table 2: Overall distribution of error types identified by Gemini 2.5 Pro as LLM-as-Judge across all 4,118 flagged reasoning mismatches.
8 Conclusion and Future Work
We introduced EngChain, a symbolic benchmark for verifiable, multi-step engineering reasoning. Built from a taxonomy of 90 problem templates across 9 domains and 3 engineering branches, EngChain provides a verifiable evaluation of multi-step engineering problem-solving, assessing the entire reasoning chain rather than only the final answer.
Our comprehensive evaluation of frontier models reveals a stark “right for the wrong reasons” phenomenon: a universal gap between modest Final Answer Accuracy and critically low Reasoning (F1 Score)s. Our qualitative analysis further reveals that this gap is not only due to error, but that a striking majority of flagged mismatches are actually Alternative Correct solutions, which single-path verifiers would wrongly penalize. For genuine failures, we find that Conceptual Errors are the dominant failure mode, not simple calculation mistakes.
Future work will proceed in several key directions. First, we will complete our planned evaluation by benchmarking the science-enhanced, math-enhanced, and general-purpose open models to provide a holistic comparison across all model categories. Second, we will expand the LLM-as-Judge analysis to other top-performing models to create diagnostic profiles of their unique failure modes. Third, to rigorously validate our qualitative methodology, we will conduct two crucial follow-up studies. A Human Evaluation Study will be performed to establish a definitive ground-truth for our judge’s classifications. This will be complemented by a Cross-Judge Consistency study to ensure our error categorizations are not an artifact of our chosen evaluator (GPT-4o (OpenAI, 2024)). This consistency study will involve a diverse set of judges, including a model from a different leading developer, such as Claude 4 Opus (Anthropic, 2025c), and a high-capability open-source model, such as Qwen 2.5 (Qwen, 2024). Finally, upon completion of these studies, all evaluation scripts, prompts, and detailed results will be publicly released to ensure full reproducibility.
Limitations
We acknowledge several limitations in this work that we plan to address in future research.
First, a significant limitation lies in our primary quantitative metric, the Reasoning (F1 Score). By design, this metric measures alignment against a single, pre-defined ground-truth path. However, our own qualitative analysis (Section 7.5) revealed that a striking majority of steps flagged as “mismatches” were actually Alternative Correct solutions. This indicates that our quantitative F1 score, when used in isolation, significantly underestimates the true reasoning capabilities of the models and inherently penalizes valid, alternative problem-solving approaches. Future work should therefore focus on developing more advanced evaluation methods, such as semantic alignment metrics or dynamic, model-based verifiers, which can programmatically accept and validate multiple correct reasoning paths.
Second, EngChain is entirely synthetic, generated from symbolic templates. While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering. More complex, ill-defined problems often require interpreting non-textual information, such as technical diagrams, schematics, or CAD models, which are not part of our current evaluation. This can be addressed in future versions by expanding the benchmark to include multi-modal tasks and by using real-world engineering documents as seeds for semi-structured problem generation.
Third, the benchmark’s scope, while covering 90 problem types, is currently limited to three core engineering branches. It does not yet include other major fields, such as civil, aerospace, or biomedical engineering. A clear next step is to expand the EngChain taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
Appendix A LLM Prompt for Domain Validation
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and rank the most fundamental, cornerstone domains for the three primary branches of engineering. The goal is to identify the non-overlapping, core subject areas that are prerequisites for almost all other advanced topics.
For example, for Chemical Engineering, ‘‘Reaction Kinetics" would be a fundamental domain, while a niche topic like ‘‘Polymer Rheology" would be a sub-specialization, not a cornerstone.
Please provide a ranked list of the top 5--7 most fundamental domains for each of the following three branches:
1. Chemical Engineering
2. Electrical Engineering
3. Mechanical Engineering
Return your answer only in a strict JSON format as given below.
{ ‘‘chemical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘electrical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘mechanical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
The above prompt was used to query large language models to cross-reference and validate the selection of fundamental engineering domains identified in our literature review.
Appendix B LLM Prompt for Area Validation
The following prompt was used to query large language models to identify the fundamental areas or sub-topics within each core engineering domain.
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and list the most fundamental, non-overlapping ‘‘areas" or ‘‘sub-topics" within the provided engineering domain.
A ‘‘fundamental area" is a major pedagogical unit or chapter within a core course on this domain. The areas you list should be distinct pillars of the subject. For example, if the domain is ‘‘Thermodynamics," a fundamental area would be ‘‘The First Law" or ‘‘Properties of Pure Fluids." A concept like ‘‘Enthalpy" would be too specific. Similarly, within a ‘‘Circuit Analysis" domain, ‘‘Kirchhoff’s Voltage Law" and ‘‘Kirchhoff’s Current Law" are foundational concepts often taught together and would be too granular to list separately.
Please provide a list of the top 3-5 most fundamental areas for the following engineering domain: [DOMAIN_NAME_HERE]
Return your answer only in a strict JSON format as given below.
{ ‘‘domain": ‘‘[DOMAIN_NAME_HERE]",
 ‘‘areas": [
   ‘‘Area 1",
   ‘‘Area 2",
   ‘‘Area 3",
   ‘‘Area 4"
Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix C Foundational Textbooks for Problem Selection
The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 3. These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngChain is pedagogically sound and grounded in the core engineering curriculum. They were also used to validate our selection of fundamental domains and areas.
Domain Textbook
Reaction Kinetics Elements of Chemical Reaction Engineering by H. Scott Fogler
Thermodynamics Introduction to Chemical Engineering Thermodynamics by J. M. Smith, H. C. Van Ness, and M. M. Abbott
Transport Phenomena Transport Phenomena by R. Byron Bird, Warren E. Stewart, and Edwin N. Lightfoot
Electromagnetics and Waves Fundamentals of Applied Electromagnetics by F. T. Ulaby and Umberto Ravaioli
Communication Systems Digital Communications by John G. Proakis and Masoud Salehi
Signals and Systems Discrete-Time Signal Processing by Alan V. Oppenheim and Ronald W. Schafer
Mechanics of Materials Mechanics of Materials by Ferdinand P. Beer, E. Russell Johnston, Jr., and John T. DeWolf
Fluid Mechanics Fundamentals of Fluid Mechanics by Bruce R. Munson, Donald F. Young, and Theodore H. Okiishi
Vibrations and Acoustics Mechanical Vibrations by Singiresu S. Rao
Table 3: Foundational textbooks used for selecting domain-specific problems.
Appendix D LLM Prompt for Pedagogical Scoring
The following prompt was used to query large language models (specifically, gemini-2.0-flash) to evaluate the pedagogical significance of each sub-topic within its parent domain.
You are an expert academic and senior curriculum designer at a top-tier engineering university with decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to evaluate the pedagogical significance of a specific engineering ‘‘area" within a broader ‘‘domain" on a scale of 1 to 5, based on the rubric below.
Rubric:
Score 5 (Cornerstone): A foundational, prerequisite topic for nearly all other concepts in the domain. Score 3 (Core Concept): A standard, important topic that builds upon cornerstone principles but is not necessarily a universal prerequisite. Score 1 (Specialized Application): A focused application or an integrative topic taught after cornerstones and core concepts are mastered. Please evaluate the following area: Domain: [DOMAIN_NAME_HERE] Area: [AREA_NAME_HERE] { ‘‘area": ‘‘[AREA_NAME_HERE]",
 ‘‘pedagogical_significance_score": <integer, a score from 1 to 5>,
 ‘‘justification": <string, a one-sentence explanation for the score>.
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix E Detailed Domain-Aware Parameterization Examples
To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngChain constrains all generated values by the principles of the domain being tested. The following provides specific examples of this methodology for each engineering branch.
E.0.1 Chemical Engineering
  • Reaction Kinetics and Stoichiometry: Instead of abstract chemicals, problems select from curated lists of common gas, liquid, and biochemical reactants like Propane, Benzene, and Glucose. More importantly, many problems are built around pre-validated, balanced chemical equations to ensure that the fundamental principles of mass conservation and stoichiometry are carefully upheld from the outset.
  • Thermodynamics: Problems involving thermodynamic calculations use a multi-layered approach to realism. For common substances, a complete and consistent set of critical properties ($T_{c},P_{c},V_{c}$) is retrieved. For reaction calorimetry, problems utilize tabulated standard heats of formation ($\Delta H_{f}^{\circ}$). Crucially, for heat transfer calculations, the model moves beyond constant specific heats by incorporating temperature-dependent heat capacity parameters ($C_{p}/R=A+BT+CT^{2}+DT^{-2}$), ensuring a high degree of physical accuracy across a range of conditions.
  • Transport Phenomena and Fluid Mechanics: To ensure realism in fluid dynamics problems, templates draw from extensive tables of fluid properties, including the density and viscosity of common liquids and gases. The benchmark also includes parameters for non-Newtonian fluids (e.g., Polymer Solutions & Melts etc), allowing it to test more complex, real-world rheological behaviors beyond simple Newtonian assumptions.
E.0.2 Electrical Engineering
  • Electromagnetics and Waves: Instead of using an arbitrary number, the phase velocity of a wave is determined by selecting a specific propagation medium from a curated list (e.g., Distilled Water, Polyethylene, Fused Silica etc). This ensures that all subsequent calculations for parameters like wavelength or wave number are physically consistent with that material’s properties. The generation process is further constrained by fundamental physical constants, including the speed of light in a vacuum ($c_{0}$) and the permittivity of free space ($\epsilon_{0}$).
  • Signals and Systems: In scenarios bridging continuous and discrete-time domains, such as sampling, the sampling frequency is intentionally set well above the Nyquist rate to model ideal conversion and isolate the effects of the discrete-time system. For purely discrete-time problems, parameters are procedurally constrained to test specific concepts. For instance, signal frequencies are generated as rational multiples of $\pi$, and downsampling problems are deliberately constructed to either cause or avoid aliasing, directly testing the understanding of frequency folding. For fundamental operations like convolution, sequences are kept short and integer-valued to focus the analysis on the computational process.
  • Digital Communications: For deterministic and random signal analysis, probabilities for discrete random variables are generated as exact rational numbers that sum perfectly to one. In digital modulation scheme problems, carrier frequencies are constrained to be integer multiples of the bit rate to simplify analysis, energy per bit ($E_{b}$) and signal-to-noise ratios ($E_{b}/N_{0}$) are generated in physically typical ranges (e.g., picojoules and realistic dB values), and modulation orders ($M$) are restricted to powers of two. Furthermore, data rates are assigned practical units such as kbps, Mbps, or Gbps, grounding the problems in contemporary engineering contexts.
E.0.3 Mechanical Engineering
  • Mechanics of Materials: Problems are grounded in the behavior of real-world materials by co-selecting properties from comprehensive tables. When a material such as “6061-T6 Aluminum” or “Carbon Fiber Reinforced Polymer” is chosen, its physically consistent Young’s Modulus ($E$), Poisson’s Ratio ($\nu$), and Shear Modulus ($G$) are used for all calculations. This ensures that problems involving stress, strain, and torsional deformation are realistic. Properties are available in both SI (GPa) and US Customary (ksi) units to reflect practical engineering work.
  • Fluid Mechanics: Realism in fluid statics problems is achieved by utilizing extensive lists of fluid and solid densities. This allows for the accurate calculation of hydrostatic pressure and buoyancy forces for a wide range of scenarios, such as a steel object submerged in seawater. The parameterization is context-aware, selecting dense fluids like Mercury for manometer problems while using lighter fluids like oil or water for pipe flow scenarios.
  • Vibrations and Acoustics: For problems involving dynamic systems, parameters such as mass, spring stiffness, and damping coefficients are randomized within ranges that represent plausible physical objects, from small mechanical components to larger structures. Additionally, the amplitudes and frequencies of harmonic forcing functions are chosen in relation to the system’s natural frequency to ensure the dynamic response is well-defined and clearly demonstrates key physical phenomena like resonance or beating.
Appendix F Authoritative Data Sources for Parameterization
To ensure the physical and engineering realism of EngChain, we built our domain-aware parameterization (discussed in Section 4.4) on authoritative data. We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 4.
Engineering Branch Authoritative Data Sources
Chemical Engineering Perry’s Chemical Engineers’ Handbook
NIST Chemistry WebBook
CRC Handbook of Chemistry and Physics
Electrical Engineering IEEE 100: The Authoritative Dictionary of IEEE Standards Terms
Standard Handbook for Electrical Engineers
ART-DEIT Database
Mechanical Engineering ASM Handbook, Volume 1
Marks’ Standard Handbook for Mechanical Engineers
Shigley’s Mechanical Engineering Design
Table 4: Primary data sources referenced for engineering parameterization across domains.
Appendix G LLM Prompt for AI-Assisted Quality Assurance (QA)
System Prompt
You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngChain benchmark.
Your task is to meticulously evaluate a new problem template based on its source code and several example outputs.
Analyze the provided information and then respond ONLY with a single, valid JSON object that strictly adheres to the schema described in the user prompt. Do not add any explanatory text or markdown formatting around the JSON object.
User Prompt
Please evaluate the following engineering problem template.
1. Template Source Code:
python {template_code}
2. Generated Instances from the Template:
Instance 1:
- Question: ‘‘{q1}"
- Solution: ‘‘{s1}"
Instance 2:
- Question: ‘‘{q2}"
- Solution: ‘‘{s2}"
Instance 3:
- Question: ‘‘{q3}"
- Solution: ‘‘{s3}"
3. Evaluation Rubric & JSON Schema:
Evaluate the template based on the rubric below. The human_review_flag should be true if any score is less than 4. The explanation should be a concise, one-sentence justification for the scores and the flag.
{ ‘‘physical_plausibility_score": <integer, a score from 1--5 based on whether the problem respects the laws of physics and engineering>,
‘‘mathematical_correctness_score": <integer, a score from 1--5 based on whether the equations and calculations are accurate>,
‘‘pedagogical_clarity_score": <integer, a score from 1--5 based on whether the problem statement is clear, unambiguous, and solvable>,
‘‘confidence_score": <integer, a score from 1--5 indicating your confidence in this evaluation>,
‘‘human_review_flag": <boolean, true if the template requires human inspection, otherwise false>,
‘‘explanation": "<string, a concise, one-sentence justification for the scores and flag>"
Appendix H Multi-Axis Rubric for LLM-Based Template Validation
As part of our AI-Assisted Quality Assurance (QA) workflow, the LLM Peer Reviewer evaluates each template against a set of criteria. The multi-axis rubric used for this validation, detailing each scoring category and its description, is defined in Table 5.
Score Category Description
Physical Plausibility Does the problem scenario respect the laws of physics and engineering?
Mathematical Correctness Are the equations used correct, and is the final calculation accurate based on the inputs?
Pedagogical Clarity Is the problem statement clear, unambiguous, and solvable with the information provided?
Confidence Score How confident is the LLM in its own validation of this template?
Human Review Flag Based on the scores above, does this template require human inspection?
Explanation A concise, one-sentence justification for the scores and the flag.
Table 5: Multi-axis rubric used for LLM-based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
Appendix I LLM-as-Judge Error Categories
During the second stage of our evaluation framework (Section 5), the LLM-as-a-Judge system performs a qualitative error diagnosis on failed reasoning steps. This provides deeper insight than a quantitative score alone by distinguishing between fundamental reasoning failures and valid alternative solutions. The five distinct error categories used for this classification are defined in Table 6.
Error Category Description
Conceptual Error The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction).
Calculation Error The model used the correct formula but made a mathematical mistake (e.g., $2\times 3=5$).
Input Error The model used the correct formula but pulled the wrong number from the problem context or a previous step.
Alternative Correct The model’s step is valid and logically sound but follows a different method or phrasing.
Other The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors.
Table 6: Error taxonomy used to categorize model mistakes during LLM-as-Judge evaluation.
Appendix J LLM Prompt for Automated Error Analysis (LLM-as-a-Judge)
You are an expert engineering professor acting as an automated evaluator. Your task is to analyze the ‘‘MODEL’S STEP" against the ‘‘GROUND-TRUTH STEP" and return a structured JSON object based on the rules and inputs below. CRITICAL RULES: 1. Your ‘‘explanation" MUST logically justify your chosen ‘‘error_category". Do not contradict yourself. For example, do not choose ‘‘Calculation Error" and then state that the calculation is correct. 2. If the model’s step is factually correct but takes a different path than the ground-truth, you MUST use the ‘‘Alternative Correct" category. Do not classify a correct step as ‘‘Other". 3. The final output must be only a raw JSON object. Do not include any introductory text, concluding remarks, or markdown formatting. Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction). ‘‘Calculation Error": The model used the correct formula but made a mathematical mistake (e.g., 2 * 3 = 5). ‘‘Input Error": The model used the correct formula but pulled the wrong number from the problem context or a previous step. ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step. ‘‘Other": The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors. Input for Analysis:
 [CONTEXT]: {problem_context}
 [GROUND-TRUTH STEP]: {gt_step}
 [MODEL’S STEP]: {pred_step}
OUTPUT FORMAT: You must now provide your analysis. Your entire response will be a single, raw JSON object. Adhere strictly to the following format with exactly two keys:
{"error_category": "...", "explanation": "..."}
Appendix K Evaluated Model Details
This appendix provides a comprehensive overview of all Large Language Models discussed in our experimental framework (Section 6). As outlined in the main text, our evaluation suite is structured around a four-way taxonomy to enable a structured comparison across different training paradigms. These four categories are: (1) Frontier Proprietary LLMs, (2) Science / Engineering Enhanced LLMs, (3) Math Enhanced LLMs, and (4) General Purpose Open LLMs.
Table 7 provides the complete list of models within each category, detailing their developing organization, parameter size (where public), backbone model (for fine-tuned variants), and source identifier (either an API endpoint or a HuggingFace repository path).
Model Organization Size Backbone Source
Frontier Proprietary LLMs
GPT-5 OpenAI N/A gpt-5-2025-08-07
GPT-4.1 OpenAI N/A gpt-4.1-2025-04-14
GPT-5 mini OpenAI N/A gpt-5-mini-2025-08-07
GPT-4.1 mini OpenAI N/A gpt-4.1-mini-2025-04-14
Claude Sonnet 4.5 Anthropic N/A claude-sonnet-4-5-20250929
Claude Sonnet 4 Anthropic N/A claude-sonnet-4-20250514
Claude Sonnet 3.7 Anthropic N/A claude-3-7-sonnet-20250219
Gemini 2.5 Pro Google DeepMind N/A gemini-2.5-pro-2025
Gemini 2.5 Flash Google DeepMind N/A gemini-2.5-flash-2025
DeepSeek V3.2 DeepSeek N/A deepseek-v3.2-20250929
DeepSeek V3.1 DeepSeek N/A deepseek-v3.1-20250922
DeepSeek R1 DeepSeek N/A deepseek-r1-20250120
Grok 4 Heavy xAI N/A grok-4-heavy-0709
Grok 4 Fast xAI N/A grok-4-fast-reasoning
Science / Engineering Enhanced LLMs
Galactica Meta AI 120B facebook/galactica-120b
Minerva Google DeepMind 540B deepmind/minerva
ChemLLM Shanghai AI Lab 13B meta-llama/Llama-2-13B ShanghaiAILab/ChemLLM-13B
SciLit01 TII 7B Qwen/Qwen2-7B-Instruct tiiuae/SciLit-01-7B
SciAgent OpenBioLLM Team 8B mistralai/Mistral-7B-v0.3 OpenBioLLM/SciAgent-8B
Math Enhanced LLMs
WizardMath WizardLM Team 7B mistralai/Mistral-7B-v0.1 WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B EleutherAI/llemma-7b meta-math/MetaMath-7B-V1.0
Mathstral Mistral AI 7B mistralai/Mistral-7B-v0.1 mistralai/Mathstral-7B-v0.1
Qwen2.5-Math Qwen Team 7B Qwen/Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
General Purpose Open LLMs
LLaMA 3.1 Meta 8B meta-llama/Llama-3.1-8B
Qwen 2.5 Alibaba Qwen Team 7B Qwen/Qwen2.5-7B-Instruct
Qwen 3 Alibaba Qwen Team 8B Qwen/Qwen3-8B
Table 7: Overview of all models evaluated, grouped by functionality: (1) frontier proprietary LLMs representing current performance upper bounds, (2) science/engineering-enhanced models trained on technical corpora, (3) math-enhanced models for quantitative reasoning, and (4) general-purpose open LLMs providing broad foundations. For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
Appendix L AI-Assisted QA Workflow: Template Examples
This appendix provides visual evidence of the LLM Peer Reviewer’s output for an automatically approved template (provided in Figure 9) and a flagged template (provided in Figure 10), illustrating the function of our automated Quality Assurance (QA) process (Section 4.5).
Figure 9: LLM Peer Review result for an Approved Template (template_rackett_equation_volume). The template, designed for Chemical Engineering Thermodynamics, receives the maximum score of 5/5 across all core categories (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity), indicating that the problem structure is robust and ready for benchmark inclusion.
Figure 10: LLM Peer Review result for a Flagged Template (template_euclidean_distance_binary). The model detected a fundamental flaw, assigning a score of 3/5 for both Plausibility and Correctness. The explicit explanation flags a “calculation error that sets sqrt(Eb) to zero,” which immediately forces the template out of the automated pipeline and into human review for manual correction.
Appendix M Detailed LLM-as-Judge Error Analysis
This appendix provides the full, fine-grained breakdown of the qualitative error analysis performed on Gemini 2.5 Pro, using GPT-4o as the judge. The following tables detail the distribution of all 4,118 reasoning steps that were flagged by the quantitative verifier as a mismatch from the ground-truth solution. Table 8 breaks down the errors by engineering branch,  Table 9 shows the distribution by problem difficulty, and Figure 11 provides a proportional breakdown of error types across all nine engineering domains.
Branch Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Chemical 1353 206 67 17 43 1686
Electrical 827 221 23 24 94 1189
Mechanical 865 207 63 40 67 1242
Total 3045 634 153 81 204 4117
Table 8: Branch-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three main engineering branches.
Difficulty Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Easy 1088 162 52 15 61 1378
Intermediate 1178 204 65 51 73 1571
Advanced 779 268 36 15 70 1168
Total 3045 634 153 81 204 4117
Table 9: Difficulty-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three problem difficulty levels.
Figure 11: Domain-level Error Analysis for Gemini 2.5 Pro. This 100% stacked bar chart illustrates the proportional distribution of error types across all nine engineering domains. This visualization highlights how the ratio of error types changes by domain; for instance, Fluid Mechanics and Signals & Systems show a much larger proportion of Conceptual Errors and Other errors compared to Thermodynamics or Reaction Kinetics, where Alternative Correct solutions dominate.