### v2

# EngTrace: A Symbolic Benchmark for Verifiable Process Supervision of Engineering Reasoning

Abstract
Large Language Models (LLMs) are increasingly entering specialized, safety-critical engineering workflows governed by strict quantitative standards and immutable physical laws, making rigorous evaluation of their reasoning capabilities imperative. However, existing benchmarks such as MMLU, MATH, and HumanEval assess isolated cognitive skills, failing to capture the physically grounded reasoning central to engineering, where scientific principles, quantitative modeling, and practical constraints must converge. To enable verifiable process supervision in engineering, we introduce EngTrace, a symbolic benchmark comprising 90 templates across three major engineering branches, nine core domains and 20 distinct areas. Through domain-aware parameterization, we generate 1,350 unique, contamination-resistant test cases to stress-test generalization. Moving beyond outcome matching, we introduce a verifiable two-stage evaluation framework that uses a tiered protocol to validate intermediate reasoning traces alongside final answers through automated procedural checks and a heterogeneous AI Tribunal. Our evaluation of 24 leading LLMs reveals a distinct trade-off between numeric precision and trace fidelity, identifying a complexity cliff where abstract mathematical pre-training fails to translate into the integrative reasoning required for advanced engineering tasks.
EngTrace: A Symbolic Benchmark
for Verifiable Process Supervision of Engineering Reasoning
Ayesha Gull1 Muhammad Usman Safder1 Rania Hossam2 Fan Zhang3 Veselin Stoyanov2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI 3The University of Tokyo Project Code App
1 Introduction
With the expansion of LLMs into high-stakes domains such as engineering, the need for their comprehensive evaluation has become paramount (Zhao et al., 2023; Xie et al., 2023a). Progress is currently gated by benchmark quality; without tests that enforce physically grounded reasoning, distinguishing genuine capability from sophisticated mimicry remains a challenge.
Figure 1: An example from EngTrace demonstrating CSTR Volume Calculation. A symbolic template generates a unique problem instance paired with a gold-standard reasoning trace to enable verifiable process supervision.
Extending the symbolic-template paradigm pioneered in mathematics (Mirzadeh et al., 2024) and recently in finance (Xie et al., 2025) to the interdependent physical systems and specialized design constraints of engineering, we introduce EngTrace, a new benchmark designed to enable verifiable process supervision of reasoning in engineering problem-solving. It is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique problem instances. Figure 1 illustrates a Continuous Stirred-Tank Reactor (CSTR) volume calculation example from our benchmark. Unlike static benchmarks, this generated instance requires synthesizing interdependent stoichiometry and reaction kinetics variables to derive the reactor size required for a specific outlet flow rate, yielding a complex reasoning trace fully verifiable against the template’s ground truth.
EngTrace addresses two critical gaps that compromise the validity of current LLM evaluation paradigms for engineering: benchmark saturation and disciplinary fragmentation. First, static benchmarks like Glue (Wang et al., 2019) and BBH (Kazemi et al., 2025) are increasingly prone to model saturation (Ott et al., 2022) and data contamination (Deng et al., 2024). EngTrace addresses these risks through symbolic templates and domain-aware parameterization, generating unique, physically grounded problems that prevent rote memorization. Second, most benchmarks evaluate critical skills in disciplinary silos. General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., Math (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively. This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012). EngTrace addresses this structural limitation through integrative reasoning templates. Unlike existing benchmarks (e.g., Zhou et al., 2025; Mudur et al., 2025) that rely solely on outcome matching, our templates demand holistic, procedural reasoning. This is a crucial requirement in engineering, where a flawed process can lead to catastrophic failure (Leveson, 2012).
In summary, our contributions are threefold: (1) We introduce EngTrace, a gold-standard symbolic benchmark for verifiable process supervision, comprising 90 problem templates spanning three major engineering branches, nine core domains, and 20 distinct areas. (2) We propose a verifiable two-stage evaluation framework that employs a tiered verification protocol utilizing automated procedural checks and a multi-model AI tribunal to assess reasoning process quality in addition to final answer accuracy. (3) We evaluate 24 models across various categories, revealing critical weaknesses in reasoning quality, with non-frontier models failing to generalize beyond basic patterns and performing poorly under complex engineering constraints.
2 Related Work
Benchmarks in Mathematics and Coding.
Extensive work has established baselines for isolated cognitive skills. In mathematics, benchmarks ranging from GSM8K (Cobbe et al., 2021) and Math (Hendrycks et al., 2021b) to HardMath (Fan et al., 2024), Putnam-Axiom (Gulati et al., 2024), and GSM-Symbolic (Mirzadeh et al., 2024) rigorously test abstract logical deduction and reasoning stability. Similarly, coding evaluations span from function-level generation in HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to real-world software engineering in SWE-Bench (Jimenez et al., 2024). While foundational, these benchmarks lack engineering utility, as proficiency in abstract logic does not translate to reliable reasoning under physical constraints (Heesch et al., 2025). Recently, FinChain (Xie et al., 2025) utilized symbolic templates to evaluate financial reasoning. While this demonstrated the utility of verifiable workflows, EngTrace extends the approach to physically grounded simulation, where reasoning traces are governed by immutable conservation laws and interdependent system dynamics.
Benchmarks in the Physical Sciences.
To bridge the gap between abstract mathematics and physical reality, recent benchmarks have pivoted to the physical sciences. While UGPhysics (Xu et al., 2025) and PhysReason (Zhang et al., 2025a) effectively test the application of reasoning to natural laws, others target advanced symbolic (Zhang et al., 2025b; Qiu et al., 2025) or qualitative (Wang et al., 2023; Shojaee et al., 2025) understanding. Despite this progress, these benchmarks remain confined to theoretical physics and lack the integrative context of engineering.
Benchmarks in Engineering.
This fragmentation extends to engineering evaluation. While generalist suites like MMLU (Hendrycks et al., 2021a),BIG-Bench (Luo et al., 2024), and SuperGPQA (Du et al., 2025) include engineering topics, their reliance on knowledge-based multiple-choice questions restricts assessment to factual recall rather than procedural fidelity. Conversely, specialized benchmarks offer depth but are constrained by narrow disciplinary scopes or a lack of scalable reasoning verification. For instance, specialized benchmarks like TransportBench, APBench, and EEE-Bench are siloed within narrow sub-disciplines (Syed et al., 2024; Chen et al., 2025; Cheng et al., 2024; Li et al., 2025; Liu and others, 2025). Others focus on tangential skills, such as design generation (Guo et al., 2025b) or software proficiency (Mudur et al., 2025), rather than fundamental reasoning, while broader efforts like EngiBench (Zhou et al., 2025) rely on subjective rubric-based scoring.
Evidently, existing benchmarks remain limited to outcome matching and fail to validate the physically grounded reasoning required for engineering. EngTrace addresses this limitation by introducing a symbolic benchmark that pairs unique problem instances with gold-standard reasoning traces to enable verifiable process supervision.
3 EngTrace Design and Methodology
3.1 Taxonomy and Content Selection
The taxonomy of EngTrace, illustrated in Figure 2, systematically covers nine core domains and 20 distinct areas across three engineering branches. This structure is not arbitrarily defined but follows a principled, two-pronged methodology to ensure it mirrors formal engineering curricula.
Figure 2: EngTrace’s Hierarchical Taxonomy spanning nine domains across three branches: process equilibrium and dynamics (Chemical), field theory and signal transmission (Electrical), and physical mechanics and material properties (Mechanical).
Domain and Area Selection.
To identify core domains and critical areas, we synthesized engineering curricular standards from ABET (ABET, 2024) and major professional societies (ASME (ASME, 2025), IEEE (IEEE, 2025), AIChE (AIChE, 2025)). We validated these foundational units by cross-referencing them with an expert-persona LLM to ensure alignment between traditional pedagogy and modern AI capabilities (prompts provided in Appendix A and Appendix B).
Template Selection.
We developed 90 symbolic templates by sourcing core engineering principles, problem typologies, and typical constraints from foundational textbooks (e.g., Fogler, 2020; Ulaby and Ravaioli, 2015), ensuring pedagogical grounding (see Appendix C for the complete list of source texts). The template distribution per area is non-uniform and guided by a Pedagogical Significance Score (1–5) derived through an expert-persona LLM (detailed prompt provided in Appendix D). This weighting ensures hierarchical fidelity: for instance, within the Reaction Kinetics domain, the core area of Mole Balances (Score: 5) includes five templates, while the specialized Levenspiel Plot Interpretation (Score: 1) is represented by a single template.
3.2 Template Generation Pipeline
The core of EngTrace is the symbolic template, a Python blueprint for generating unique, physically grounded engineering problems. As shown in Figure 3, the pipeline first samples domain-aware parameters from authoritative data sources (Appendix E) and fundamental principles (see detailed examples in Appendix F). It then performs core computations with multi-layer sanity checks to derive computationally verified ground-truth traces. These are serialized into natural language (question, solution) pairs, creating a dynamic problem space that effectively resists data contamination.
3.3 Template Validation
To establish EngTrace as a high-fidelity “Gold Standard” benchmark, we utilize a robust two-stage validation process that pairs automated heuristic auditing with expert manual verification.
Stage 1: Automated AI Pre-Screening.
Building upon the “Panel of LLM” (PoLL) methodology (Verga et al., 2024), we construct an “AI Tribunal” comprising three frontier reasoning models: GPT-5 (OpenAI, 2025a), Claude Opus 4.5 (Anthropic, 2025b), and Gemini 3 (Google, 2025). Each model acts as an independent judge, utilizing the specific prompt provided in Appendix G, to evaluate the source code and generated instances against the strict multi-axis rubric detailed in Appendix H.
To mitigate potential hallucinations and ensure output reliability, we apply a hybrid aggregation strategy requiring each template $T$ to satisfy three consensus conditions. First, the Median Quality Score ($S_{med}$) across the tribunal must reach a threshold $\theta=4$ for every evaluative dimension $d\in D$:
$\forall d\in D,\quad S_{med}^{(d)}=\text{median}(S_{m_{1}}^{(d)},S_{m_{2}}^{(d)},S_{m_{3}}^{(d)})\geq 4$ (1)
Second, we enforce a Majority Vote safety check, discarding any template where more than one judge identifies a violation ($\sum F_{i}\geq 2$, where $F\in\{0,1\}$).
Finally, we quantify uncertainty using a Disagreement Score ($\sigma_{max}$), defined as the maximum standard deviation across all dimensions. Templates exhibiting high variance are flagged for manual review to resolve potential logical ambiguities:
$\sigma_{max}=\max_{d\in D}\left(\sqrt{\frac{1}{3}\sum_{i=1}^{3}(S_{m_{i}}^{(d)}-\mu^{(d)})^{2}}\right)\leq 0.5$ (2)
where $\mu^{(d)}$ represents the mean score across judges for dimension $d$.
Stage 2: Expert Human Certification.
Templates passing the AI tribunal undergo a double-blind review by domain experts through a custom Streamlit interface presented in Appendix I. We assess Inter-Annotator Agreement per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility (Approve/Reject) and Krippendorff’s Alpha ($\alpha$) for ordinal quality ratings (1–5), the latter accounting for disagreement magnitude and missing data.
AI-Human Alignment.
To assess validation reliability, we measure correspondence between the AI Tribunal and human experts using False Positive Rate (FPR), to quantify filter safety (templates approved by AI but rejected by human experts), and Spearman’s Rank Correlation ($\rho$) to evaluate monotonic ranking consistency, chosen over Pearson to better handle the ordinal nature and high-quality skew of the data. Detailed results are provided in Appendix J.
Figure 3: The EngTrace Template Generation Pipeline maps domain-aware constraints to serialized question-solution pairs via physically-validated computation, illustrated here using a CSTR volume calculation example.
3.4 Difficulty Scaling
Problem difficulty in EngTrace is calibrated through a multi-axis framework defined by three dimensions: Conceptual Complexity (isolated principles vs. multi-domain synthesis), Mathematical Sophistication (direct algebraic substitution vs. differential equations), and Procedural Depth (length of interdependent reasoning chains). A detailed quantitative breakdown of the dataset’s difficulty distribution is provided in Appendix K.
4 Evaluation Framework
Inspired by recent advances in process supervision (Lightman et al., 2023) and cascaded model verification (Chen et al., 2023), we introduce a two-stage evaluation framework that validates intermediate reasoning steps in addition to final-answer accuracy through a tiered verification protocol. To ensure alignment with professional standards, we validated this protocol through a blind human-expert study, confirming that our Tribunal-based approach significantly outperforms standard metrics in capturing engineering reasoning quality (see Appendix M for full methodology and ablation results).
4.1 Preliminaries
We define the evaluation of a reasoning chain as a comparison between the ground truth solution $G$ and the model-generated predicted solution $P$. Both are represented as ordered sequences of $M$ and $N$ discrete reasoning steps, respectively:
$G=(g_{1},g_{2},\dots,g_{M}),\quad P=(p_{1},p_{2},\dots,p_{N})$ (3)
Each individual reasoning step $s_{i}\in\{G,P\}$ consists of a textual derivation $t_{i}$ and, where applicable, an associated numerical or symbolic intermediate result $a_{i}=\text{Res}(s_{i})$.
To assess the validity of the derivation, we seek a mapping that aligns steps in $P$ to their logical counterparts in $G$. Unlike standard string matching, this alignment must account for structural variations such as step merging, reordering, or alternative valid methodologies while ensuring that the cumulative logic remains faithful to the underlying engineering principles.
4.2 Final Answer Accuracy
To evaluate the correctness of the final predicted outcome, let $g_{M}$ and $p_{N}$ denote the last steps of the ground truth and predicted solutions, from which we extract the terminal numerical results $A_{G}=\text{Res}(g_{M})$ and $A_{P}=\text{Res}(p_{N})$. A response is deemed correct if the relative error between these values satisfies the following condition:
$\mathbb{I}_{final}=\mathbb{I}\left(\frac{|A_{G}-A_{P}|}{|A_{G}|}<\epsilon_{final}\right)$ (4)
where $\epsilon_{final}=0.02$ (2%), a strict threshold required to prevent catastrophic error propagation in high-precision engineering derivations.
4.3 Textual Quality
We evaluate the fluency and semantic fidelity of generated solutions using standard reference-based metrics (Xie et al., 2023b). We employ BERTScore (Zhang et al., 2020) to measure deep semantic similarity via contextual embedding alignment, and ROUGE-2 (Lin, 2004) to quantify lexical overlap. Together, these metrics assess the model’s ability to employ appropriate domain-specific terminology.
4.4 Tiered Reasoning Verification
To evaluate the full reasoning chain, we characterize the validity of the predicted sequence $P$ against the ground truth $G$ through a tiered verification protocol. This process consists of two phases: determining step-level validity through hierarchical assessment and aggregating these results into a global reasoning metric.
Step-Level Validity Assessment.
We compute a validity matrix $V\in\mathbb{R}^{M\times N}$, where each element $V_{i,j}$ represents the validity of predicted step $p_{j}$ relative to ground truth step $g_{i}$. This validity is determined by two successive tiers of verification.
Tier 1: Automated Symbolic Verification.
This tier fills the matrix $V$ by enforcing syntactic and numerical precision through a binary step-validity score:
$S_{\mathcal{T}_{1}}(g_{i},p_{j})=\delta_{num}\land\delta_{sem}$ (5)
where $S_{\mathcal{T}_{1}}=1$ only if the numerical consistency (relative error tolerance $\epsilon=0.02$) and semantic similarity (Cross-Encoder threshold $\tau=0.7$) are both satisfied. To optimize computational efficiency, we calculate the alignment ratio, which represents the proportion of steps in $P$ successfully verified by Tier 1. If this ratio exceeds a confidence threshold ($\rho=0.8$) and the final answer is correct, the reasoning chain is classified as a standard derivation and accepted without further review. Chains falling below this threshold are escalated to Tier 2 for heuristic assessment.
Tier 2: Multi-Model Heuristic Verification.
Escalated chains are evaluated by a heterogeneous AI Tribunal to identify valid alternative derivations (prompt provided in Appendix L). Each judge $J_{k}$ classifies the observed discrepancy between $p_{j}$ and $g_{i}$ into a category $C_{k}$, which is then mapped to a scalar score:
$\mu(C)=\begin{cases}1.0&\text{if }C=\text{Alternate Correct}\\ 0.5&\text{if }C=\text{Calculation Error}\\ 0.0&\text{if }C\in\{\text{Conceptual Error, Other}\}\end{cases}$ (6)
The majority consensus of these scalar scores defines $S_{\mathcal{T}_{2}}$, which is assigned to $V_{i,j}$ to finalize the validity matrix for non-standard derivations. In the absence of a clear majority, the assessment defaults to a conservative tie-breaker where the lowest score is assigned.
Global Reasoning Metric $\mathbf{F1_{rec}}$.
To quantify the overall quality of the reasoning chain, we calculate the Recovered Reasoning F1 Score ($F1_{rec}$). We first identify the optimal alignment that maximizes the aggregate verification score:
$\mathcal{R}_{total}=\max_{\sigma\in\mathcal{A}}\sum_{(i,j)\in\sigma}V_{i,j}$ (7)
where $\mathcal{A}$ represents the set of valid one-to-one mappings between steps. This optimization is solved using the Hungarian matching algorithm on the finalized validity matrix $V$. We define recovered precision and recall as $P_{rec}=\mathcal{R}_{total}/N$ and $R_{rec}=\mathcal{R}_{total}/M$, respectively, where $N$ and $M$ are the number of predicted and ground truth steps. Finally, $F1_{rec}$ is computed as the harmonic mean of $P_{rec}$ and $R_{rec}$.
5 Experiments
5.1 Evaluation Model Suite
We evaluate 24 models categorized into a three-way taxonomy to isolate the impacts of scale, access, and specialization: (1) Frontier proprietary models represent the current state-of-the-art in reasoning and instruction following. Our evaluation includes 12 models in this category: GPT-5 {Base, Mini} and GPT-4.1 {Base, Mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025c, d, a), Gemini {3 Pro, 2.5 Pro, 2.5 Flash} (Google, 2025; Comanici et al., 2025), and DeepSeek {V3, R1} (Liu et al., 2024; Guo et al., 2025a). (2) General-purpose open models provide strong, domain-agnostic foundations and allow for transparent analysis of performance relative to parameter count. We evaluate 8 models: Meta-Llama-3.1 {8B, 70B} (Grattafiori et al., 2024), Qwen2.5 {7B, 14B, 72B} and Qwen3-8B (Qwen, 2024, 2025), and Gemma {2-9B, 3-27B} (Gemma Team et al., 2024; Gemma Team and Google DeepMind, 2025). (3) Math-enhanced models are systems specifically trained on mathematical and symbolic corpora to enhance quantitative reasoning. We evaluate 4 specialized models in the 7B parameter class: Qwen2.5-Math-7B (Yang et al., 2024), Mathstral-7B (Mistral, 2024), WizardMath-7B (Luo et al., 2023), and MetaMath-7B (Yu et al., 2023). Detailed configurations and model sources can be found in Appendix N.
5.2 Experimental Setup
We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model. We evaluate all models under a unified decoding configuration: a low temperature of 0.2 to ensure deterministic outputs and a maximum token limit of 4,096 to allow for detailed reasoning. All models are evaluated in a zero-shot setting using a standardized prompt and output parsing pipeline; full implementation specifications are provided in Appendix O.
5.3 Results
5.3.1 Overall Model Performance
Table 1 summarizes performance across five key metrics: Final Answer Accuracy (FAC), Reasoning F1, and standard semantic measures (BERTScore, ROUGE-2/L). The results reveal a distinct hierarchy: frontier proprietary models define the state-of-the-art, with Gemini 3 Pro achieving peak numerical precision (FAC 64.00%) and DeepSeek R1 leading in logical process alignment (Reasoning F1 44.41%). In contrast, open-weights models lag significantly in both metrics despite competitive semantic scores, even among large-scale systems like Llama 3.1 70B. Notably, specialized math-enhanced models in the 7B parameter class generally underperform general-purpose counterparts of equivalent size, suggesting that abstract mathematical pre-training does not strictly transfer to the physically constrained reasoning required for engineering.
Model Final Answer $\uparrow$ Reasoning (F1) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
Frontier Proprietary LLMs
Claude 3.7 Sonnet 57.63 39.36 88.09 29.42 37.36
Claude 4 Sonnet 56.81 36.80 87.92 29.52 37.22
Claude 4.5 Sonnet 51.78 33.61 87.37 28.06 35.39
DeepSeek V3 62.37 37.06 86.54 25.69 34.57
DeepSeek R1 61.04 44.41 86.48 25.58 34.49
Gemini 2.5 Flash 59.78 42.22 87.64 29.14 36.70
Gemini 2.5 Pro 60.44 39.38 88.24 31.88 38.07
Gemini 3 Pro 64.00 37.05 86.71 27.32 34.99
GPT-4.1 57.11 31.22 86.63 26.78 34.46
GPT-4.1 Mini 61.56 36.87 86.50 25.06 33.87
GPT-5 61.04 41.81 88.19 28.98 36.55
GPT-5 Mini 63.70 41.01 87.95 26.63 34.74
General Purpose Open LLMs
Gemma 2 9B 26.59 21.37 88.19 28.02 35.77
Gemma 3 27B 44.30 27.59 87.04 23.93 32.66
Llama 3.1 8B 23.70 18.61 87.65 25.59 32.50
Llama 3.1 70B 40.81 30.04 88.34 28.87 35.82
Qwen 2.5 7B 35.63 23.67 86.30 24.21 32.89
Qwen 2.5 14B 42.07 27.67 86.44 24.94 33.38
Qwen 2.5 72B 48.07 28.02 86.46 25.58 33.96
Qwen 3 8B 23.70 16.68 86.91 8.93 11.09
Math Enhanced LLMs
Mathstral 7B 19.70 18.59 87.79 28.02 36.42
MetaMath 7B 6.07 2.37 82.97 10.57 16.69
Qwen 2.5 Math 7B 23.04 18.18 85.61 21.71 29.98
WizardMath 7B 8.22 6.38 83.66 14.14 20.38
Table 1: Overall Performance Comparison of 24 models on the EngTrace benchmark. Scores are reported as percentages. The Final Answer column denotes strict numerical accuracy, while Reasoning (F1) measures the structural alignment of the derivation steps. The highest score in each column is highlighted in bold.
5.3.2 Performance by Engineering Branch
Analysis by branch reveals a trade-off between reasoning fidelity and numerical precision. Chemical Engineering proved the most rigorous, yielding the lowest average scores; DeepSeek R1 achieved a decisive lead here (FAC 64.89%, F1 40.91%), confirming that complex stoichiometry demands sustained chain-of-thought capabilities. In Electrical Engineering, Gemini 3 Pro secured the top FAC (68.67%), while DeepSeek R1 retained the highest Reasoning F1 (41.22%), indicating structurally superior derivations even when numeric outputs diverge. Mechanical Engineering was the most accessible, with Gemini 2.5 Pro achieving the benchmark’s highest overall FAC (71.56%). Notably, Llama 3.1 70B showed strong semantic performance here (BERTScore 88.37%), suggesting mechanical terminology is particularly well-represented in general pre-training corpora.
5.3.3 Performance by Engineering Domain
A granular analysis of specific domains (Figure 4) reveals critical variances in capability. In Chemical Engineering, DeepSeek R1 dominates Thermodynamics and Reaction Kinetics (FAC 78.33% and 58.00%), suggesting its reinforcement learning optimization effectively navigates coupled equations; conversely, Transport Phenomena remains a bottleneck (max 53.33% FAC). Electrical Engineering results are heterogeneous: Gemini 3 Pro excels in Digital Communications (77.33% FAC), likely due to its logic-based nature, but physics-heavy domains like Electromagnetics remain challenging, with no model exceeding 56.00% FAC. In Mechanical Engineering, Mechanics of Materials appears largely solved (Gemini 3 Pro: 92.67% FAC), while Fluid Mechanics caps at 56.00% FAC. This indicates that while models have mastered static relationships, they continue to struggle with the non-linear differential equations governing fluid dynamics.
Figure 4: Domain-Level Performance comparison of the top four models ranked by mean Final Answer Accuracy. The plot reveals a “spiky” performance profile, highlighting that model capabilities are highly specialized and not generalized across topics.
5.3.4 Performance by Difficulty Level
Stratifying performance by difficulty (Figure 5) reveals a functional transition in model utility. In the Easy tier, efficiency-oriented architectures like GPT-5 Mini (67.84%) and open models like Llama 3.1 70B (52.55%) perform considerably well, suggesting flagship compute is unnecessary for fundamental engineering problems. However, the Advanced tier exposes a sharp divergence: Llama 3.1 70B collapses to 23.33%, identifying frontier systems as the sole viable option for high-complexity scenarios, led by Gemini 3 Pro (54.55%). Notably, specialized math-enhanced models (e.g., Qwen 2.5 Math) consistently underperform, peaking at just 32.35% on Easy tasks, confirming that abstract mathematical pre-training is insufficient for domain-specific physical reasoning.
Figure 5: Difficulty-Level Performance of the top four models by mean Final Answer Accuracy. The plot shows a “complexity cliff”: efficiency-oriented models degrade sharply at advanced tasks, whereas frontier reasoning models remain stable.
5.4 Error Analysis
To characterize reasoning failures beyond simple accuracy metrics, we analyzed the 17,528 reasoning chains (54.1% of the dataset) that failed Tier 1 verification. This audit reveals that across all models, 85.86% of traces flagged as discrepancies were identified by the AI Tribunal as Alternative Correct derivations, as illustrated in Figure 6. This rate approached saturation in frontier models (97.17%), confirming that advanced models possess a high degree of solution diversity that renders rigid template-matching metrics unsuitable for engineering evaluation. Focusing on confirmed failures, the aggregate data shows that Conceptual Errors (1,318) outnumbered Calculation Errors (711) by a ratio of nearly 2:1. This indicates that while models frequently hallucinate physics principles, they are comparatively more reliable in arithmetic calculations. This trend correlates strongly with complexity: while Easy tasks exhibited a balanced error distribution, Advanced tasks saw the conceptual-to-calculation failure ratio widen to 3.7:1. Finally, 795 “Lucky Guesses” (2.5%), defined as correct answers with structurally invalid reasoning, underscore the necessity of multi-stage verification over final-answer accuracy alone.
Figure 6: Distribution of Identified Reasoning Errors across model categories. While Frontier models primarily utilize valid alternative methods, General Open and Math-Enhanced models exhibit a significantly higher density of confirmed reasoning failures.
6 Conclusion and Future Work
In this work, we introduced EngTrace, a symbolic benchmark designed to enable verifiable process supervision of physically grounded engineering reasoning in LLMs. Our evaluation of 24 models reveals a “complexity cliff,” where efficiency-oriented and open-weights models degrade steeply in advanced engineering tasks while frontier reasoning models maintain stability. These findings highlight a critical reasoning gap, emphasizing that trustworthy engineering AI requires more than just general logic. Future work will explore tool-augmented reasoning to decouple semantic physical logic from computational precision and investigate uncertainty quantification to determine if models can reliably calibrate confidence in safety-critical engineering scenarios.
Limitations
We acknowledge several limitations in this work that we plan to address in future research.
First, our current evaluation treats models as isolated inference engines, detached from the practical constraints of applied engineering. By focusing solely on static text generation, we overlook two critical components of reliable design: iterative self-correction and risk assessment. Consequently, while we effectively measure raw theoretical proficiency, our current framework does not fully capture the stability or safety assurance required for deployment in high-stakes engineering workflows.
Second, EngTrace is entirely synthetic, generated from symbolic templates. While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering. More complex, ill-defined problems often require interpreting non-textual information, such as technical diagrams, schematics, or CAD models, which are not part of our current evaluation. This can be addressed in future versions by expanding the benchmark to include multi-modal tasks and by using real-world engineering documents as seeds for semi-structured problem generation.
Third, the benchmark’s scope, while covering 90 problem types, is currently limited to three core engineering branches. It does not yet include other major fields, such as civil, aerospace, or biomedical engineering. A clear next step is to expand the EngTrace taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
Ethical Statement and Broad Impact
This work utilizes synthetic data generated through symbolic templates and standard engineering principles; no private, sensitive, or proprietary industrial data was used. While EngTrace aims to advance reliable engineering reasoning, we emphasize that current LLMs should not be deployed in safety-critical physical systems without rigorous, hardware-in-the-loop verification. We hope this benchmark fosters the development of interpretable, verifiable, and physically grounded AI for high-stakes engineering applications.
Data License
The EngTrace dataset and associated code will be released under the MIT License.
Acknowledgments
We extend our gratitude to the engineering professionals and domain experts for validating the EngTrace framework and reviewing the templates.
Appendix A LLM Prompt for Domain Validation
To identify core domains, we consulted a panel of four frontier models: GPT-4o, Claude 3.5 Sonnet, Gemini 2.5 Pro, and DeepSeek V3. Each was prompted individually to identify cornerstone engineering domains. We finalized the nine core domains via majority voting. The domain identification prompt is provided below.
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and rank the most fundamental, cornerstone domains for the three primary branches of engineering. The goal is to identify the non-overlapping, core subject areas that are prerequisites for almost all other advanced topics.
For example, for Chemical Engineering, ‘‘Reaction Kinetics" would be a fundamental domain, while a niche topic like ‘‘Polymer Rheology" would be a sub-specialization, not a cornerstone.
Please provide a ranked list of the top 5--7 most fundamental domains for each of the following three branches:
1. Chemical Engineering
2. Electrical Engineering
3. Mechanical Engineering
Return your answer only in a strict JSON format as given below.
{ ‘‘chemical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘electrical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘mechanical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix B LLM Prompt for Area Validation
Following the domain validation procedure, we utilized the same model panel and aggregated the pedagogical areas via majority voting. The prompt used for area identification is provided below.
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and list the most fundamental, non-overlapping ‘‘areas" or ‘‘sub-topics" within the provided engineering domain.
A ‘‘fundamental area" is a major pedagogical unit or chapter within a core course on this domain. The areas you list should be distinct pillars of the subject. For example, if the domain is ‘‘Thermodynamics," a fundamental area would be ‘‘The First Law" or ‘‘Properties of Pure Fluids." A concept like ‘‘Enthalpy" would be too specific. Similarly, within a ‘‘Circuit Analysis" domain, ‘‘Kirchhoff’s Voltage Law" and ‘‘Kirchhoff’s Current Law" are foundational concepts often taught together and would be too granular to list separately.
Please provide a list of the top 3-5 most fundamental areas for the following engineering domain: [DOMAIN_NAME_HERE]
Return your answer only in a strict JSON format as given below.
{ ‘‘domain": ‘‘[DOMAIN_NAME_HERE]",
 ‘‘areas": [
   ‘‘Area 1",
   ‘‘Area 2",
   ‘‘Area 3",
   ‘‘Area 4"
Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix C Foundational Textbooks for Template Selection
The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 2. These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngTrace is pedagogically sound and grounded in the core engineering curriculum. They were also used to validate our selection of fundamental domains and areas.
Domain Textbook
Reaction Kinetics Elements of Chemical Reaction Engineering by H. Scott Fogler
Thermodynamics Introduction to Chemical Engineering Thermodynamics by J. M. Smith, H. C. Van Ness, and M. M. Abbott
Transport Phenomena Transport Phenomena by R. Byron Bird, Warren E. Stewart, and Edwin N. Lightfoot
Electromagnetics and Waves Fundamentals of Applied Electromagnetics by F. T. Ulaby and Umberto Ravaioli
Communication Systems Digital Communications by John G. Proakis and Masoud Salehi
Signals and Systems Discrete-Time Signal Processing by Alan V. Oppenheim and Ronald W. Schafer
Mechanics of Materials Mechanics of Materials by Ferdinand P. Beer, E. Russell Johnston, Jr., and John T. DeWolf
Fluid Mechanics Fundamentals of Fluid Mechanics by Bruce R. Munson, Donald F. Young, and Theodore H. Okiishi
Vibrations and Acoustics Mechanical Vibrations by Singiresu S. Rao
Table 2: Foundational textbooks used for selecting domain-specific problems.
Appendix D LLM Prompt for Pedagogical Scoring
The following prompt was used to query LLMs (described in Section A) to evaluate the pedagogical significance of each sub-topic within its parent domain.
You are an expert academic and senior curriculum designer at a top-tier engineering university with decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to evaluate the pedagogical significance of a specific engineering ‘‘area" within a broader ‘‘domain" on a scale of 1 to 5, based on the rubric below.
Rubric:
Score 5 (Cornerstone): A foundational, prerequisite topic for nearly all other concepts in the domain. Score 3 (Core Concept): A standard, important topic that builds upon cornerstone principles but is not necessarily a universal prerequisite. Score 1 (Specialized Application): A focused application or an integrative topic taught after cornerstones and core concepts are mastered. Please evaluate the following area: Domain: [DOMAIN_NAME_HERE] Area: [AREA_NAME_HERE] { ‘‘area": ‘‘[AREA_NAME_HERE]",
 ‘‘pedagogical_significance_score": <integer, a score from 1 to 5>,
 ‘‘justification": <string, a one-sentence explanation for the score>.
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix E Authoritative Data Sources for Parameterization
To ensure the physical and engineering realism of EngTrace, we built our domain-aware parameterization on authoritative data. We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 3.
Engineering Branch Authoritative Data Sources
Chemical Engineering Perry’s Chemical Engineers’ Handbook
NIST Chemistry WebBook
CRC Handbook of Chemistry and Physics
Electrical Engineering IEEE 100: The Authoritative Dictionary of IEEE Standards Terms
Standard Handbook for Electrical Engineers
ART-DEIT Database
Mechanical Engineering ASM Handbook, Volume 1
Marks’ Standard Handbook for Mechanical Engineers
Shigley’s Mechanical Engineering Design
Table 3: Primary data sources referenced for engineering parameterization across domains.
Appendix F Detailed Domain-Aware Parameterization Examples
To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngTrace constrains all generated values by the principles of the domain being tested. The following provides specific examples of this methodology for each engineering branch.
F.0.1 Chemical Engineering
  • Reaction Kinetics and Stoichiometry: Instead of abstract chemicals, problems select from curated lists of common gas, liquid, and biochemical reactants like Propane, Benzene, and Glucose. More importantly, many problems are built around pre-validated, balanced chemical equations to ensure that the fundamental principles of mass conservation and stoichiometry are carefully upheld from the outset.
  • Thermodynamics: Problems involving thermodynamic calculations use a multi-layered approach to realism. For common substances, a complete and consistent set of critical properties ($T_{c},P_{c},V_{c}$) is retrieved. For reaction calorimetry, problems utilize tabulated standard heats of formation ($\Delta H_{f}^{\circ}$). Crucially, for heat transfer calculations, the model moves beyond constant specific heats by incorporating temperature-dependent heat capacity parameters ($C_{p}/R=A+BT+CT^{2}+DT^{-2}$), ensuring a high degree of physical accuracy across a range of conditions.
  • Transport Phenomena and Fluid Mechanics: To ensure realism in fluid dynamics problems, templates draw from extensive tables of fluid properties, including the density and viscosity of common liquids and gases. The benchmark also includes parameters for non-Newtonian fluids (e.g., Polymer Solutions & Melts etc), allowing it to test more complex, real-world rheological behaviors beyond simple Newtonian assumptions.
F.0.2 Electrical Engineering
  • Electromagnetics and Waves: Instead of using an arbitrary number, the phase velocity of a wave is determined by selecting a specific propagation medium from a curated list (e.g., Distilled Water, Polyethylene, Fused Silica etc). This ensures that all subsequent calculations for parameters like wavelength or wave number are physically consistent with that material’s properties. The generation process is further constrained by fundamental physical constants, including the speed of light in a vacuum ($c_{0}$) and the permittivity of free space ($\epsilon_{0}$).
  • Signals and Systems: In scenarios bridging continuous and discrete-time domains, such as sampling, the sampling frequency is intentionally set well above the Nyquist rate to model ideal conversion and isolate the effects of the discrete-time system. For purely discrete-time problems, parameters are procedurally constrained to test specific concepts. For instance, signal frequencies are generated as rational multiples of $\pi$, and downsampling problems are deliberately constructed to either cause or avoid aliasing, directly testing the understanding of frequency folding. For fundamental operations like convolution, sequences are kept short and integer-valued to focus the analysis on the computational process.
  • Digital Communications: For deterministic and random signal analysis, probabilities for discrete random variables are generated as exact rational numbers that sum perfectly to one. In digital modulation scheme problems, carrier frequencies are constrained to be integer multiples of the bit rate to simplify analysis, energy per bit ($E_{b}$) and signal-to-noise ratios ($E_{b}/N_{0}$) are generated in physically typical ranges (e.g., picojoules and realistic dB values), and modulation orders ($M$) are restricted to powers of two. Furthermore, data rates are assigned practical units such as kbps, Mbps, or Gbps, grounding the problems in contemporary engineering contexts.
F.0.3 Mechanical Engineering
  • Mechanics of Materials: Problems are grounded in the behavior of real-world materials by co-selecting properties from comprehensive tables. When a material such as “6061-T6 Aluminum” or “Carbon Fiber Reinforced Polymer” is chosen, its physically consistent Young’s Modulus ($E$), Poisson’s Ratio ($\nu$), and Shear Modulus ($G$) are used for all calculations. This ensures that problems involving stress, strain, and torsional deformation are realistic. Properties are available in both SI (GPa) and US Customary (ksi) units to reflect practical engineering work.
  • Fluid Mechanics: Realism in fluid statics problems is achieved by utilizing extensive lists of fluid and solid densities. This allows for the accurate calculation of hydrostatic pressure and buoyancy forces for a wide range of scenarios, such as a steel object submerged in seawater. The parameterization is context-aware, selecting dense fluids like Mercury for manometer problems while using lighter fluids like oil or water for pipe flow scenarios.
  • Vibrations and Acoustics: For problems involving dynamic systems, parameters such as mass, spring stiffness, and damping coefficients are randomized within ranges that represent plausible physical objects, from small mechanical components to larger structures. Additionally, the amplitudes and frequencies of harmonic forcing functions are chosen in relation to the system’s natural frequency to ensure the dynamic response is well-defined and clearly demonstrates key physical phenomena like resonance or beating.
Appendix G LLM Prompt for Automated AI Pre-Screening
You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngTrace benchmark.
Your task is to meticulously evaluate a new problem template based on its source code and several example outputs.
Analyze the provided information and then respond ONLY with a single, valid JSON object that strictly adheres to the schema described in the user prompt. Do not add any explanatory text or markdown formatting around the JSON object.
Please evaluate the following engineering problem template.
1. Template Source Code:
python {template_code}
2. Generated Instances from the Template:
Instance 1:
- Question: ‘‘{q1}"
- Solution: ‘‘{s1}"
Instance 2:
- Question: ‘‘{q2}"
- Solution: ‘‘{s2}"
Instance 3:
- Question: ‘‘{q3}"
- Solution: ‘‘{s3}"
3. Evaluation Rubric & JSON Schema:
Evaluate the template based on the rubric below. The human_review_flag should be true if any score is less than 4. The explanation should be a concise, one-sentence justification for the scores and the flag.
{ ‘‘physical_plausibility_score": <integer, a score from 1--5 based on whether the problem respects the laws of physics and engineering>,
‘‘mathematical_correctness_score": <integer, a score from 1--5 based on whether the equations and calculations are accurate>,
‘‘pedagogical_clarity_score": <integer, a score from 1--5 based on whether the problem statement is clear, unambiguous, and solvable>,
‘‘confidence_score": <integer, a score from 1--5 indicating your confidence in this evaluation>,
‘‘human_review_flag": <boolean, true if the template requires human inspection, otherwise false>,
‘‘explanation": "<string, a concise, one-sentence justification for the scores and flag>"
Appendix H Multi-Axis Rubric for Automated AI Pre-Screening
During automated AI pre-screening, judges in the AI Tribunal evaluate templates against a set of criteria. The evaluation rubric, detailing scoring categories and descriptions, is provided in Table 4.
Score Category Description
Physical Plausibility Does the problem scenario respect the laws of physics and engineering?
Mathematical Correctness Are the equations used correct, and is the final calculation accurate based on the inputs?
Pedagogical Clarity Is the problem statement clear, unambiguous, and solvable with the information provided?
Confidence Score How confident is the LLM in its own validation of this template?
Human Review Flag Based on the scores above, does this template require human inspection?
Explanation A concise, one-sentence justification for the scores and the flag.
Table 4: Multi-axis rubric used for AI Tribunal based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
Appendix I Human Certification Interface
To facilitate the rigorous Stage 2 evaluation, we developed a custom web-based annotation tool using the Streamlit framework. This interface, illustrated in Figure 7, was designed to streamline the expert review process while ensuring data consistency for statistical analysis.
Figure 7: The EngTrace Expert Annotation Interface. The dashboard features a dual-view layout for verifying source code against rendered traces, enabling experts to input structured quality scores and definitive admissibility decisions.
The workflow within the interface proceeds as follows:
  1. 1.
    Dual-View Inspection: The main panel displays the rendered natural language Question and step-by-step Solution Trace exactly as they appear in the final dataset. Simultaneously, reviewers can toggle a tab to inspect the underlying Python Source Code, allowing them to verify the algorithmic logic alongside the output.
  2. 2.
    Dynamic Re-sampling: To verify consistency beyond a single static snapshot, the interface allows reviewers to interactively regenerate instances. This enables experts to stress-test the parameterization logic and confirm that physical constraints hold across multiple random seeds.
  3. 3.
    Ordinal Quality Scoring: Reviewers evaluate the template across three dimensions (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity) using interactive sliders (1–5). These values provide the ordinal data required for calculating Krippendorff’s Alpha ($\alpha$).
  4. 4.
    Binary Certification Gate: The “Certification” toggle forces a definitive Approve or Reject decision, generating the categorical data for Fleiss’ Kappa ($\kappa$).
  5. 5.
    Mandatory Feedback Loop: If a template is rejected, the interface enforces a mandatory text input field for specific feedback (e.g., "The mole balance assumption is invalid for variable volume"). This feedback is logged and routed back to the authors for immediate correction.
Appendix J Template Validation Results
J.1 Stage 1: AI Tribunal Impact
In the initial audit, the AI Tribunal identified 26 Critical Failures and 6 Controversial templates. These templates were manually refactored based on the specific failure modes identified by the models (e.g., unit mismatches or unrealistic material properties) until a 100% pass rate was achieved; representative examples of these identified failures and their associated reasoning are provided in Table 5.
Template Metadata Median Score Votes Model Model Feedback Summary
adiabatic_flame_temperature Chemical Engineering Thermodynamics 2 2/3 GPT-5 Structure clear but temperatures too low; likely $C_{p}$ polynomial or unit mismatch in energy balance.
Claude 4.5 Opus Method sound but values physically inaccurate (e.g., Acetylene $\sim$1493K vs $\sim$2500K); suspects data parameter issues.
Gemini 3 Pro Critical unit error identified in heat capacity integration (incorrect application of gas constant $R$).
multi_segment_rod Mechanical Engineering Mechanics of Materials 3 2/3 GPT-5 Equations correct but ignores yield limits, resulting in unrealistically high stresses.
Claude 4.5 Opus Physics generally sound, but questions the validity of applying large tensile loads to concrete.
Gemini 3 Pro Identifies physically impossible scenarios where brittle materials are subjected to loads exceeding fracture limits.
Table 5: Qualitative Feedback from the AI Tribunal. Examples of critical failures identified during initial screening, categorized by model-specific technical critiques.
J.2 Stage 2: Inter-Annotator Agreement
We measured agreement among domain experts per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility and Krippendorff’s Alpha ($\alpha$) for ordinal quality metrics. As shown in Table 6, we observed perfect agreement ($\kappa=1.0$) across all domains regarding the final inclusion of templates. This unusually high consensus suggests that the AI Tribunal based stage 1 successfully acted as a rigorous filter, ensuring that only high-quality candidates reached the human review stage.
Domain Templates Fleiss’ $\kappa$ Interpretation
Chem. Eng. 30 1.00 Perfect Agreement
Elec. Eng. 30 1.00 Perfect Agreement
Mech. Eng. 30 1.00 Perfect Agreement
Table 6: Binary Admissibility Consensus (Fleiss’ Kappa). Agreement on whether a template should be Included/Rejected.
Krippendorff’s Alpha ($\alpha$)
Domain
Physical
Plausibility
Math.
Correctness
Pedagogical
Clarity
Chem. Eng. 0.224 1.00 0.457
Elec. Eng. 0.372 1.00 0.332
Mech. Eng. 0.647 1.00 0.078
GLOBAL 0.437 1.00 0.269
Table 7: Ordinal Quality Consistency (Krippendorff’s Alpha). Agreement on 1–5 scoring across three axes.
For the ordinal quality ratings, Table 7 reveals a clear distinction between objective and subjective metrics. Mathematical Correctness achieved perfect agreement ($\alpha=1.0$), reflecting the objective nature of verifying derivations. Physical Plausibility showed moderate agreement ($\alpha\approx 0.44$), as experts occasionally debated edge cases in parameter realism. Conversely, Pedagogical Clarity exhibited lower agreement ($\alpha\approx 0.27$), highlighting the inherent subjectivity in evaluating educational "style."
J.3 AI-Human Alignment
We evaluated the performance of the AI Tribunal (Stage 1) against the Human Experts (Stage 2) using two key metrics:
  1. 1.
    Filter Safety (False Positive Rate): The AI Tribunal demonstrated perfect safety (FPR $0.00\%$), as zero templates approved by the AI were subsequently rejected by humans (Target $10\%$ achieved).
  2. 2.
    Scoring Alignment (Spearman’s Rho): The calculated Spearman’s Rho was low ($\rho=0.065,p=0.54$), indicating weak rank correlation. This is likely an artifact of range restriction: since the AI filter was highly effective, the surviving templates all clustered near the maximum quality score (4–5/5). With variance effectively eliminated (a "ceiling effect"), rank correlation becomes statistical noise rather than a meaningful signal of misalignment.
Appendix K Dataset Statistics
Structural Balance and Granularity.
Each branch is represented by exactly 30 templates, ensuring that no single discipline dominates the aggregate performance metrics. This uniform top-level distribution controls for branch-specific bias. At the granular level, the number of Areas (6–7 per branch) reflects the specific curricular density of each discipline; for instance, Mechanical Engineering consolidates multiple sub-topics into broad areas like Fluid Mechanics, resulting in slightly fewer total areas (6) compared to Chemical and Electrical (7), without sacrificing concept coverage.
Difficulty Distribution.
The dataset difficulty is intentionally stratified to mirror curricular progression. We prioritize foundational and applied reasoning (Level 1 & 2: $\approx$75%) to test core engineering laws and procedural workflows, while reserving the most complex synthesis tasks for the advanced tier (Level 3: $\approx$25%). This weighting ensures models are evaluated primarily on standard practice while still distinguishing elite capabilities in multi-physics integration. Table 8 presents the detailed statistical breakdown.
Branch Structure Difficulty Level Total
Dom. Areas Easy Int. Adv.
Chemical 3 7 13 9 8 30
Electrical 3 7 11 12 7 30
Mechanical 3 6 10 13 7 30
Total 9 20 34 34 22 90
Table 8: Dataset Statistics by Branch. The dataset features a perfectly balanced distribution of 30 templates per branch, with difficulty levels scaled to test both foundational knowledge and advanced synthesis.
Appendix L LLM Prompt for Multi-Model Heuristic Verification
You are an expert engineering professor acting as an automated evaluator. Your task is to analyze the ‘‘MODEL’S STEP" against the ‘‘GROUND-TRUTH STEP" and return a structured JSON object based on the rules and inputs below. CRITICAL RULES: 1. Your ‘‘explanation" MUST logically justify your chosen ‘‘error_category". Do not contradict yourself. For example, do not choose ‘‘Calculation Error" and then state that the calculation is correct. 2. If the model’s step is factually correct but takes a different path than the ground-truth, you MUST use the ‘‘Alternative Correct" category. Do not classify a correct step as ‘‘Other". 3. The final output must be only a raw JSON object. Do not include any introductory text, concluding remarks, or markdown formatting. Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula. ‘‘Calculation Error": The model used the correct formula but made a mathematical mistake. ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step. ‘‘Other": The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors. Input for Analysis:
 [CONTEXT]: {problem_context}
 [GROUND-TRUTH STEP]: {gt_step}
 [MODEL’S STEP]: {pred_step}
OUTPUT FORMAT: You must now provide your analysis. Your entire response will be a single, raw JSON object. Adhere strictly to the following format with exactly two keys:
{"error_category": "...", "explanation": "..."}
Appendix M Evaluation Framework Validation
To ensure that the EngTrace evaluation framework reflects expert engineering judgment, we conducted a rigorous human-alignment study. We established a validation set of 20 instantiated questions, randomly sampled with stratification across all three branches and difficulty levels, and generated responses from five representative models: deepseek-reasoner, gemini-3-pro-preview, gemma-3-27b-it, qwen2.5-math-7b, and Meta-Llama-3.1-8B-Instruct.
This procedure resulted in 100 randomized and anonymized responses. Three subject matter experts per branch independently scored them on a 1–5 Likert scale. Our initial analysis revealed a strong correlation between human-rated reasoning quality and final answer accuracy (Spearman’s $\rho=0.9130$, $p<0.05$), confirming that coherent reasoning traces are highly predictive of numerical correctness. Consequently, we prioritized reasoning process quality as the primary dimension for optimizing our automated evaluator.
M.1 Expert Evaluation Rubrics
The following rubrics were utilized by subject matter experts to evaluate the 100 validation responses. Experts provided independent scores for reasoning and accuracy on a 1–5 Likert scale as detailed in Table 9 and Table 10.
Score Rating Description
5 Excellent Flawless derivation; follows all physical laws and engineering principles.
4 Good Correct reasoning with minor notation inconsistencies or suboptimal paths.
3 Fair Generally sound logic; minor errors in parameter grounding or arithmetic.
2 Poor Significant logical gaps or incorrect governing equations for the domain.
1 Deficient Nonsensical, physically impossible, or complete hallucination of concepts.
Table 9: Expert rubric used for assessing the quality of the reasoning process.
Score Rating Description
5 Excellent Numerically correct within tolerance including correct units.
4 Good Correct value but missing or incorrectly formatted units.
3 Fair Off by a small margin due to rounding; correct magnitude.
2 Poor Numerically incorrect by a significant margin or order of magnitude.
1 Deficient Entirely wrong or absent.
Table 10: Expert rubric for assessing final answer accuracy.
M.2 Ablation Study Results
We systematically assessed various evaluator configurations to optimize alignment with these expert ratings. Table 11 shows the Spearman’s Rank Correlation ($\rho$) between automated pipeline scores and human judgment.
The results demonstrate that the combination of an AI Tribunal ($\tau=0.8$) and numeric step-matching yields the highest alignment ($\rho=0.632$). This alignment significantly degrades upon the removal of the Tribunal ($\rho=0.411$), confirming its necessity for capturing valid alternative derivations. Furthermore, the poor performance of standard NLP metrics ($\rho=0.325$) reinforces the need for the domain-specific numerical verification employed in EngTrace.
Technical Configuration Alignment ($\rho$)
Tribunal ($\tau=0.8$) + Numeric Step-Matching 0.632
Tribunal ($\tau=0.95$) + Numeric Step-Matching 0.581
Tribunal ($\tau=0.6$) + Numeric Step-Matching 0.525
No AI Tribunal; Strict Ground-Truth Matching 0.411
Tribunal + Step-Matching (Numbers ignored) 0.469
ROUGE-L / BERTScore (Standard NLP metrics) 0.325
Table 11: Ablation results showing the Spearman’s rank correlation ($\rho$) between automated reasoning scores and human expert ratings. The data highlights the necessity of the AI Tribunal and numerical verification.
Appendix N Evaluated Model Details
Table 12 details the 24 models evaluated in our experiments (Section 5). We report the developing organization, parameter size, backbone architecture (where applicable), and the specific access identifier used to ensure reproducibility.
Model Organization Size Backbone Source
Frontier Proprietary LLMs
GPT-5 Base OpenAI N/A gpt-5-base-2025-08-07
GPT-5 Mini OpenAI N/A gpt-5-mini-2025-08-07
GPT-4.1 Base OpenAI N/A gpt-4.1-2025-04-14
GPT-4.1 Mini OpenAI N/A gpt-4.1-mini-2025-04-14
Claude Sonnet 4.5 Anthropic N/A claude-sonnet-4-5-20250929
Claude Sonnet 4 Anthropic N/A claude-sonnet-4-20250514
Claude Sonnet 3.7 Anthropic N/A claude-3-7-sonnet-20250219
Gemini 3 Pro Google N/A gemini-3-pro-001
Gemini 2.5 Pro Google N/A gemini-2.5-pro-002
Gemini 2.5 Flash Google N/A gemini-2.5-flash-002
DeepSeek V3 DeepSeek N/A deepseek-v3
DeepSeek R1 DeepSeek N/A deepseek-reasoner
General Purpose Open LLMs
Llama 3.1 70B Meta 70B meta-llama/Llama-3.1-70B-Instruct
Llama 3.1 8B Meta 8B meta-llama/Llama-3.1-8B-Instruct
Qwen 2.5 72B Alibaba Cloud 72B Qwen/Qwen2.5-72B-Instruct
Qwen 2.5 14B Alibaba Cloud 14B Qwen/Qwen2.5-14B-Instruct
Qwen 2.5 7B Alibaba Cloud 7B Qwen/Qwen2.5-7B-Instruct
Qwen 3 8B Alibaba Cloud 8B Qwen/Qwen3-8B-Instruct
Gemma 3 27B Google 27B google/gemma-3-27b-it
Gemma 2 9B Google 9B google/gemma-2-9b-it
Math Enhanced LLMs
Qwen2.5-Math Qwen Team 7B Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
Mathstral Mistral AI 7B Mistral-7B mistralai/Mathstral-7B-v0.1
WizardMath WizardLM Team 7B Mistral-7B WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B Llemma-7B meta-math/MetaMath-7B-V1.0
Table 12: Overview of the 24 models evaluated, grouped into (1) frontier proprietary, (2) general-purpose open, and (3) math-enhanced categories. For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
Appendix O Inference Prompt and Output Processing
O.1 Standardized Inference Prompt
To ensure consistent and parsable outputs across all models, we employed a standardized zero-shot system prompt. This prompt explicitly instructs the model to structure its response with specific headings and to format the final result for automated extraction.
You are an expert engineer. Solve the following
problem by providing a detailed, structured
solution. Use the exact headings and
formatting provided below.
# Given
List all known variables and their
values with units.
# Find
State the variable(s) to be calculated.
# Formulae
Write down all necessary governing equations
before substituting any values.
# Solution
Provide a step-by-step calculation. Each step
must start on a new line and be formatted exactly
as ’**Step X:**’, where X is the step number.
Show the substitution of values into
the formulae clearly.
# Final Answer
State the final numerical result with its units
in the format: **Answer:** [value] [units]
O.2 Output Parsing Logic
To process the model outputs generated by the prompt above, we employ a heuristic parsing pipeline:
  1. 1.
    Step Segmentation: We use regular expressions to identify line-start markers matching either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.). A subsequent regex cleans these prefixes to isolate the reasoning text.
  2. 2.
    Value Extraction: We extract numerical values from each step and the final answer using a three-priority heuristic:
    • Priority 1: Matching the explicit **Answer:** tag defined in the prompt.
    • Priority 2: Identifying the last number following an equals sign (e.g., V = 100.0).
    • Priority 3: Extracting the last standalone number in the segment.
  3. 3.
    Fallback: If no overall **Answer:** tag is found, the final numerical value extracted from the last reasoning step is used for verification.